### v1

# Course Allocation with Credits via Stable Matching

Abstract
In the Course Allocation problem, there are a set of students and a set of courses at a given university. University courses may have different numbers of credits, typically related to different numbers of learning hours, and there may be other constraints such as courses running concurrently. Our goal is to allocate the students to the courses such that the resulting matching is stable, which means that no student and course(s) have an incentive to break away from the matching and become assigned to one another. We study several definitions of stability and for each we give a mixture of polynomial-time algorithms and hardness results for problems involving verifying the stability of a matching, finding a stable matching or determining that none exists, and finding a maximum size stable matching. We also study variants of the problem with master lists of students, and lower quotas on the number of students allocated to a course, establishing additional complexity results in these settings.
1 Introduction
1.1 Background and motivation
Every year, undergraduate students must select a list of courses to enrol on for the next academic year. Their selections are constrained by several factors, mainly that university courses have a limited capacity and that universities enforce a limit on the academic load of students. This load is usually expressed in terms of the total number of course credits, with longer courses that have more associated learning hours typically having more credits than shorter ones.
[…]
In some cases, all courses within a faculty have the same preferences over students elte_regulations_2025 . For example, all courses may give priority to students with higher grades over those with lower grades. We show that, when such a uniform ranking of students exists, known as a master list, it is always possible to find a stable matching in polynomial time for all four stability criteria. However, stable matchings can have different sizes, and we show that the problem of finding a stable matching of maximum size is NP-hard, except in the case of pair stability. Finally, we consider the case of courses having lower quotas biro_college_2010 , a minimum number of students that a course needs in order to run. In general we find that problems involving lower quotas are NP-hard.
1.2 Related work
The Course Allocation problem has been extensively studied in the past, with papers in this area studying a range of different models of the problem. One of the best known mechanisms for addressing this problem is the Approximate Competitive Equilibrium from Equal Incomes mechanism (A-CEEI) budish_combinatorial_2011 , in which students rank possible schedules (each schedule is a subset of courses which does not violate any constraint) and each is given an approximately equal budget to bid on courses. The outcome is approximately Pareto efficient and satisfies some other fairness criteria. It was later refined, resulting in the Course Match mechanism budish_course_2017 . Due to the difficulty in expressing an ordering across all possible schedules, later work has modified Course Match to use machine learning to better estimate the true preferences of students soumalias_machine_2024 .
Alternative approaches to the Course Allocation problem involve students ranking courses individually, instead of ranking bundles, but ensuring that the final allocation respects constraints such as the maximum number of credits a student can take or disallowing excluded combinations such as courses that run concurrently biswas_2023_algorithmic . Other types of constraints have also been considered, for example where courses have prerequisites, corequisites or similar constraints that force students, if they desire to enrol in a given course, to also enrol in a different one cechlarova_pareto_2018 .
[…]
Stable matching problems were introduced by Gale and Shapley gale_college_1962 . They considered the Stable Marriage problem (sm), a problem in which agents are divided into men and women and must be paired together. They show that a stable matching, a matching in which no man-woman pair wants to break with their partners and be together instead, always exists and can be found in polynomial time. The algorithm that they presented is known as the Deferred Acceptance mechanism (DA). The Hospitals/Residents problem (hr) is a variation of sm in which resident doctors apply to hospitals, and hospitals may have more than one resident. As in the case of sm, a stable matching always exists and can be found in polynomial time using DA gale_college_1962 . However, a variant known as Hospitals/Residents with Couples (hrc), in which some residents form a couple and submit a joint preference list, is NP-hard ng_1988_complexity ; ron_1990_np , even if each member of every couple has the same preference list as their partner mcdermid_keeping_2010 . This variant is known as Hospitals/Residents with Sizes, because each couple can be treated as a resident of size $2$ mcdermid_keeping_2010 . In the case of Course Allocation, if every course has the same number of credits, DA can be used to find a stable matching. However, as in the case of couples in hrc, having courses with different numbers of credits means that DA generally does not output a stable matching.
[…]
In many matching markets, including some Course Allocation settings elte_regulations_2025 , agents in one side are ordered according to a master list, such that the preferences of the agents on the other side are consistent with the master list. The presence of a master list guarantees the existence of a stable matching in problems such as Stable Roommates abraham_stable_2007 and makes some NP-hard problems solvable in polynomial time abraham_stable_2007 . However, this is not always the case. Generally, problems such as finding a maximum size stable matching are NP-hard even in the presence of master lists, if there are ties present abraham_stable_2007 ; irving_stable_2008 .
In some applications, courses have lower quotas, representing a minimum number of students that need to be enroled in the course in order for the course to run cechlarova_pareto_2017 , but this extension has not received a great deal of attention in the literature on Course Allocation. However, there are many results for problems such as Hospitals/Residents (sometimes in the equivalent setting of university admissions). Some papers study the case where every hospital must satisfy its lower quota; this case is known as the no closures model. If hospitals have the choice of satisfying the lower quota or being closed and not admitting any residents, the model is the closures model. It is possible to determine in polynomial time whether there exists a stable matching in the hr with lower quotas and no closures model, because in every stable matching each hospital has the same number of residents gale_1985_some ; roth_1984_evolution ; roth_1986_allocation , so one simply needs to check whether in a stable matching all hospitals satisfy their lower quotas. On the other hand, if there are ties in the preference lists of the participants, hr with lower quotas and no closures is NP-complete, because determining the existence of a complete stable matching is NP-complete iwama_2002_stable ; manlove_2002_hard , so setting lower quotas to be equal to upper quotas proves the result for hr with lower quotas and no closures. On the other hand, for hr with lower quotas and closures, the problem of deciding whether a stable matching exists is NP-complete. biro_college_2009 ; biro_college_2010 .
Our model of the Course Allocation can be placed in the context of the wider literature on matching problems with preferences. For example, if courses are assumed to not have preferences over students, then we obtain a model similar to some previously existing Course Allocation models biswas_2023_algorithmic ; budish_course_2017 , or to Capacitated House Allocation manlove_algorithmics_2013 . If both students and courses have preferences, but students can be matched with at most one course, we obtain the Hospitals/Residents problem gale_college_1962 ; mcdermid_keeping_2010 . If on the other hand courses can be matched with at most one student, every course has one or two credits, and neither students nor courses have preferences, we obtain the Matching with Couples model biro_matching_2014 . In this model finding a maximum size matching is NP-hard biro_matching_2014 .
Aziz et al. aziz_stability_2018 analyse a model similar to ours in the context of allocating refugee families to host localities. Some of their stability definitions are similar to ours; we will elaborate on this point later. Another model of interest is due to Hoyer and Stroh-Maraun in the context of the School Choice problem hoyer_stability_2020 . In their model some students may need more resources than others, which is similar to the fact that, in our model, some courses have a higher number of credits than others. They present a (not necessarily polynomial-time) algorithm that finds a stable matching if one exists. Another paper studying the same problem uses a variation of the Top-Trading Cycles algorithm to output a Pareto optimal matching stroh-maraun_weighted_2024 .
1.3 Our contributions
In this paper we introduce a new model for the Course Allocation problem, focused on courses with different numbers of credits and preferences over students, in which the goal is to find a matching that is stable, or a stable matching that is of maximum size. We also consider these problems in the presence of lower quotas. Table 1 presents a summary of our results. Each problem can be identified as follows. Course allocation under pair, first-coalition, coalition and pair-size stability is identified as ca-p, ca-fc, ca-c and ca-ps respectively. For the problem of testing whether a given matching is stable, we append -test. The problem of finding a stable matching or reporting that none exists has the suffix -find. In the presence of a master list of students, we insert -ml into the suffix. The problem of finding a maximum size stable matching is represented by -max. In the presence of lower quotas, there are two variants: either (i) every course must be assigned at least as many students as its lower quota, in which case there are no course closures (represented by -lq-nc) or (ii) a course can be closed (and not run) if it does not meet its lower quota (represented by lq-cl). The central results in this paper are those in the -find, -ml-find and -ml-max rows.
pair first-coalition coalition pair-size
ca-p ca-fc ca-c ca-ps
Testing (-test) P [O2] co-NP-c [T1] co-NP-c [T1] co-NP-c [T1]
Finding (-find) NP-h, $\not\exists$ [T3] NP-h, $\not\exists$ [T4] Open P, $\exists$ [T2]
Finding, master list P, $\exists$ [P2] P, $\exists$ [P2] P, $\exists$ [P2] P, $\exists$ [P2]
(-ml-find)
Max size, master list P [O3] NP-h [T7] NP-h [T5] NP-h [T8]
(-ml-max)
Lower quotas, master list, P, $\not\exists$ [O6] NP-h, $\not\exists$ [O4] NP-h, $\not\exists$ [O4] NP-h, $\not\exists$ [O4]
no closures (-ml-lq-nc-find)
Lower quotas, master list, NP-h, $\exists$ [T9] NP-h, $\exists$ [O5] NP-h, $\exists$ [O5] NP-h, $\exists$ [O5]
closures (-ml-lq-cl-max)
Table 1: Summary of our contributions. The rows represent the different problems of interest, while the columns refer to the different definitions of stability. The symbol $\exists$ indicates that a matching of this type always exists, while $\not\exists$ indicates that a matching of this type does not always exist. The results included in this paper are referred to by a letter, indicating whether the result is an observation (O), proposition (P) or theorem (T), followed by the observation, proposition or theorem number.
1.4 Structure of the paper
The remainder of this paper is structured as follows. Section 2 defines the problem model formally. Section 3 presents results on verifying the stability of a given matching, while Section 4 contains results related to finding a stable matching (or reporting that none exists). Section 5 covers results on finding a maximum size stable matching, and Section 6 summarises the results and presents open problems.
2 Model
In this section we begin by defining an instance of the Course Allocation (ca) problem, which will be our base problem model. An instance $I$ of ca involves a set $S$ of students and a set $C$ of courses. Each student $s_{i}\in S$ has a preference list denoted by $P(s_{i})\subseteq C$. Similarly, each course $c_{j}\in C$ has a preference list denoted by $P(c_{j})\subseteq S$. We will abuse notation by referring to preference lists as both a set and an ordered list. If $c_{j}$ appears in $s_{i}$’s preference list, we say that $c_{j}$ is acceptable to $s_{i}$, and vice versa. We assume that a student is acceptable to a course if and only if that course is acceptable to the student. Hence, we say that a student-course pair $(s_{i},c_{j})$ is acceptable if $s_{i}$ and $c_{j}$ are acceptable to each other. Given two courses $c_{j},c_{k}\in P(s_{i})$, we write $c_{j}\succ_{s_{i}}c_{k}$ if $s_{i}$ prefers $c_{j}$ to $c_{k}$. We define $\succ_{c_{j}}$ analogously for all $c_{j}\in C$.
A special case of ca called Course Allocation with Master List (ca-ml) occurs when there is a uniform ranking $\succ_{ML}$ of all students, known as a master list, where $s_{i}\succ_{ML}s_{j}$ if $s_{i}$ is placed higher than $s_{j}$ in $ML$. In the presence of a master list, for all $c_{j}\in C$, $c_{j}$ respects the master list in the sense that, for any two students $s_{i},s_{k}\in P(c_{j})$, $s_{i}\succ_{c_{j}}s_{k}$ if and only if $s_{i}\succ_{ML}s_{k}$. Throughout the paper, we assume that both individual preference lists and master lists are strict.
We assume that each course $c_{j}\in C$ has an associated number of credits, denoted by $O(c_{j})$. We extend this notation to a subset $D\subseteq C$ of courses, i.e., $O(D)=\sum_{c_{j}\in D}O(c_{j})$ is the sum of the number of credits of all courses in $D$. Also course $c_{j}$ has an upper quota (or capacity), denoted by $q^{+}(c_{j})$, which gives the maximum number of students that can be assigned to $c_{j}$. Each student $s_{i}$ has a maximum number of credits, denoted $T(s_{i})$, that she can be assigned from the courses she is allocated.
We now define the concept of a matching in instances of ca.
Definition 1.
Given an instance of ca, an assignment $M$ is a subset of acceptable student/course pairs. For a given student $s_{i}\in S$, we let $C_{M}(s_{i})=\{c_{j}\in C:(s_{i},c_{j})\in M\}$ be the set of courses that $s_{i}$ is assigned in $M$, and analogously, given a course $c_{j}\in C$, we let $S_{M}(c_{j})=\{s_{i}\in S:(s_{i},c_{j})\in M\}$ be the set of students assigned to $c_{j}$ in $M$. The assignment $M$ is a matching if additionally $M$ satisfies the following conditions:
  • for each student $s_{i}\in S$, $O(C_{M}(s_{i}))\leq T(s_{i})$, that is, $s_{i}$ cannot take more credits than her credit limit;
  • for each course $c_{j}\in C$, $|S_{M}(c_{j})|\leq q^{+}(c_{j})$, that is, $c_{j}$ cannot enrol more students than its upper quota.
Although in this work we are mostly concerned with a setting where students take any combination of acceptable courses that does not exceed their credit limit, in real life there are other factors that affect the feasibility of a matching. For example, two courses may run at the same time, students may have credit limits per quarter or semester, or at most one out of three courses from a give group may be selected, because their content is very similar. In general, such constraints can be modelled by defining the feasible subsets of courses for each student. An important special case of such subsets are downward-feasible subsets, as defined by Utture et al. utture_student_2019 which we now define here.
Definition 2 (utture_student_2019 ).
For each $s_{i}\in S$, a set of downward-feasible constraints is a subset $F_{i}$ of the power set of $P(s_{i})$ such that, if $D\in F_{i}$, then for any $D^{\prime}\subseteq D$, $D^{\prime}\in F_{i}$. A matching $M$ is feasible if, for every $s_{i}$, $C_{M}(s_{i})\in F_{i}$. Given a matching $M$ and a student $s_{i}$, we say that a course $c_{j}$ is feasible for $s_{i}$ if $(s_{i},c_{j})$ is acceptable and the matching $M\cup(s_{i},c_{j})$ is feasible.
That is, if it is feasible for a student to be matched with a set of courses, then it is feasible to be matched with a subset of these courses. From the definition it is also clear that credit limits of students are a type of downward-feasible constraints, but we will treat credit limits separately. This is because credit limits are more central to the ca problem; almost all universities and ca models have credit limits on students, while other downward-feasible constraints do not always apply. Every hardness result in this paper assumes that there are no downward-feasible constraints, and so they still apply if these are present.
Throughout this paper we will assume that we can check in polynomial time whether a given matching is feasible, and similarly whether a et of feasible constraints is downward-feasible. In general downward-feasible constraints can model many rules that can be represented efficiently (e.g., a student must take at most $60$ credits per semester, or no two courses scheduled at the same time can be assigned, etc.)
We define the size of a matching to be the total number of credits taken by all students, or alternatively, the number of students enroled in courses multiplied by their number of credits. That is, the size of a matching $M$ is $\sum_{s_{i}\in S}\sum_{c_{j}\in C_{M}(s_{i})}O(c_{j})=\sum_{c_{j}\in C}O(c_{j})\cdot|S_{M}(c_{j})|$. We say that a matching $M$ is course-complete if every course is filled to its capacity, that is, for all $c_{j}\in C$, $|S_{M}(c_{j})|=q^{+}(c_{j})$. A matching is student-complete if every student takes as many credits as possible, that is, for all $s_{i}\in S$, $O(C_{M}(s_{i}))=T(s_{i})$. The following observation is then immediate.
Observation 1.
Each of a course-complete and a student-complete matching is a maximum size matching.
Our interest lies in finding a stable matching. Informally a stable matching ensures that there is no student and course (or set of courses) that have an incentive to deviate by becoming matched together, possibly disregarding one or more of their existing assignments.
Because courses have different numbers of credits, there are several possible definitions of stability. The most natural definition is the following, which is analogous to the standard definition of stability in the Stable Marriage or Hospitals/Residents. It is equivalent to the definitions of stability in Hospitals/Residents with Sizes mcdermid_keeping_2010 and in School Choice papers by Hoyer and Stroh-Maraun hoyer_stability_2020 and Stroh-Maraun stroh-maraun_weighted_2024 . It is also similar to the definition of stability due to Aziz et al. aziz_stability_2018 .
Definition 3.
In a matching $M$, a student $s_{i}$ and a course $c_{j}$ form a blocking pair if $(s_{i},c_{j})$ is acceptable, $(s_{i},c_{j})\notin M$, and the following conditions hold:
  1. 1.
    $c_{j}$ has fewer students assigned to it than its upper quota $q^{+}(c_{j})$ or $c_{j}$ prefers $s_{i}$ over some $s_{k}$ assigned to $c_{j}$. That is, $|S_{M}(c_{j})|<q^{+}(c_{j})$ or $s_{i}\succ_{c_{j}}s_{k}$ for some $s_{k}\in S_{M}(c_{j})$.
  2. 2.
    There exists a subset $D_{i}\subseteq C_{M}(s_{i})$ (possibly empty) such that $s_{i}$ prefers $c_{j}$ to every $c_{k}\in D_{i}$, and $O(C_{M}(s_{i}))-O(D_{i})+O(c_{j})\leq T(s_{i})$. That is, $s_{i}$ has enough credits remaining for $c_{j}$, possibly after removing a subset of courses that $s_{i}$ prefers less than $c_{j}$.
  3. 3.
    If there are downward-feasible constraints, the matching $(M\cup\{(s_{i},c_{j})\})\setminus\{(s_{i},c_{k}):c_{k}\in D_{i}\}$ is feasible. That is, removing $D_{i}$ and adding $c_{j}$ to $C_{M}(s_{i})$ does not result in an infeasible matching.
If $M$ has no blocking pair then it is pair-stable. Course Allocation with pair stability is denoted ca-p.
An example of a matching with a blocking pair is given in Figure 1. However, we can see that in this example $s_{1}$ would have fewer credits if she was matched with $c_{1}$ instead of $c_{3}$. In such a situation, it may be argued that such a student is not better off, since students generally need to have a curriculum that is as complete as possible. That motivates our next definition of stability. This definition is similar to the definition of stability by demand of Aziz et al. aziz_stability_2018 .
$s_{i}$ $T(s_{i})$ $\succ_{s_{i}}$ $c_{j}$ $O(c_{j})$ $q^{+}(c_{j})$ $\succ_{c_{j}}$
[…]
$s_{2}$ 1 $\boxed{c_{2}}$ $\succ$ $c_{1}$ $c_{2}$ 1 1 $s_{1}$ $\succ$ $\boxed{s_{2}}$
$c_{3}$ 2 1 $\boxed{s_{1}}$
Figure 1: Example of a matching in a Course Allocation instance, highlighted in green, and a blocking pair, highlighted in red.
Definition 4.
In a matching $M$, a student $s_{i}$ and a course $c_{j}$ form a size-blocking pair if $(s_{i},c_{j})$ is acceptable, $(s_{i},c_{j})\notin M$, and the following conditions hold:
  1. 1.
    $c_{j}$ has fewer students assigned to it than its upper quota $q^{+}(c_{j})$ or $c_{j}$ prefers $s_{i}$ over some $s_{k}$ assigned to $c_{j}$. That is, $|S_{M}(c_{j})|<q^{+}(c_{j})$ or $s_{i}\succ_{c_{j}}s_{k}$ for some $s_{k}\in S_{M}(c_{j})$.
  2. 2.
    There exists a subset $D_{i}\subseteq C_{M}(s_{i})$ (possibly empty) such that $s_{i}$ prefers $c_{j}$ to every $c_{k}\in D_{i}$, $O(c_{j})\geq O(D_{i})$, and $O(C_{M}(s_{i}))-O(D_{i})+O(c_{j})\leq T(s_{i})$. That is, $s_{i}$ has enough credits remaining for $c_{j}$, possibly after removing a subset of courses that $s_{i}$ prefers less than $c_{j}$ and with a total number of credits smaller than or equal to that of $c_{j}$.
  3. 3.
    If there are downward-feasible constraints, the matching $(M\cup\{(s_{i},c_{j})\})\setminus\{(s_{i},c_{k}):c_{k}\in D_{i}\}$ is feasible. That is, removing $D_{i}$ and adding $c_{j}$ to $C_{M}(s_{i})$ does not result in an infeasible matching.
If $M$ has no size-blocking pair then it is pair-size-stable. Course Allocation with pair-size stability is denoted ca-ps.
The only difference between pair stability and pair-size stability is that we require that a student in a size-blocking pair must not end up with fewer credits than they started with when participating in a size-blocking pair. However, this may have unintended consequences. For example, consider agent $s_{1}$ from Figure 2. If $s_{1}$ is matched with $c_{3}$, then she would prefer to be matched with $c_{1}$ and $c_{2}$ instead. However, because $O(c_{3})=2$ and $O(c_{1})=O(c_{2})=1$, $s_{1}$ cannot form a size-blocking pair with either $c_{1}$ or $c_{2}$. This motivates our next stability definition, in which we allow students to block with more than one course.
$s_{i}$ $T(s_{i})$ $\succ_{s_{i}}$ $c_{j}$ $O(c_{j})$ $q^{+}(c_{j})$ $\succ_{c_{j}}$
[…]
$c_{3}$ 2 1 $\boxed{s_{1}}$
Figure 2: Example of a pair-stable matching in a Course Allocation instance, highlighted in green. A subset of courses that $s_{1}$ may prefer over $c_{3}$ is shown in red.
Definition 5.
In a matching $M$, a student $s_{i}$ and a set of courses $B\subseteq C$ form a blocking coalition if for all $c_{j}\in B$, $(s_{i},c_{j})$ is acceptable, for all $c_{j}\in B$, $(s_{i},c_{j})\notin M$, and the following conditions hold:
  1. 1.
    for each $c_{j}\in B$, either $c_{j}$ has fewer students assigned to it than its upper quota $q^{+}(c_{j})$ or $c_{j}$ prefers $s_{i}$ over some $s_{k}$ assigned to $c_{j}$. That is, for every $c_{j}\in B$, either $|S_{M}(c_{j})|<q^{+}(c_{j})$ or $s_{i}\succ_{c_{j}}s_{k}$ for some $s_{k}\in S_{M}(c_{j})$.
  2. 2.
    There exists a subset $D_{i}\subseteq C_{M}(s_{i})$ (possibly empty) such that $s_{i}$ prefers every $c_{j}\in B$ to every $c_{k}\in D_{i}$, $O(B)\geq O(D_{i})$ and $O(C_{M}(s_{i}))-O(D_{i})+O(B)\leq T(s_{i})$. That is, $s_{i}$ has enough credits remaining for all courses in $B$, possibly after removing a subset of courses that $s_{i}$ prefers less than all courses in $B$ and with a total number of credits smaller than or equal to that of $B$.
  3. 3.
    If there are downward-feasible constraints, the matching $(M\cup\{(s_{i},c_{j})|c_{j}\in B\})\setminus\{(s_{i},c_{k}):c_{k}\in D_{i}\}$ is feasible. That is, removing $D_{i}$ and adding $c_{j}$ to $C_{M}(s_{i})$ for every $c_{j}\in B$ does not result in an infeasible matching.
If $M$ has no blocking coalition then it is coalition-stable. Course Allocation with coalition stability is denoted ca-c.
Under coalition stability, $s_{1}$ from Figure 2 can form a blocking coalition with $c_{1}$ and $c_{2}$ if she is matched with $c_{3}$. Moving on, we now consider at Figure 3. Here, if $s_{1}$ is matched with $c_{3}$, then she cannot form a blocking coalition with $c_{1}$ and $c_{2}$, because she prefers $c_{3}$ over $c_{2}$. However, it is plausible that some students would rather be matched with $c_{1}$ and $c_{2}$ instead, because their best course in the new matching is strictly better than the best course in the old matching. Hence, we obtain a new definition of stability.
$s_{i}$ $T(s_{i})$ $\succ_{s_{i}}$ $c_{j}$ $O(c_{j})$ $q^{+}(c_{j})$ $\succ_{c_{j}}$
[…]
$c_{3}$ 2 1 $\boxed{s_{1}}$
Figure 3: Example of a coalition-stable matching in a Course Allocation instance, highlighted in green. A subset of courses that $s_{1}$ may prefer over $c_{3}$ is shown in red.
Definition 6.
In a matching $M$, a student $s_{i}$ and a set of courses $B\subseteq C$ form a first-blocking coalition if for all $c_{j}\in B$, $(s_{i},c_{j})$ is acceptable, for all $c_{j}\in B$, $(s_{i},c_{j})\notin M$, and the following conditions hold:
  1. 1.
    for each $c_{j}\in B$, either $c_{j}$ has fewer students assigned to it than its upper quota $q^{+}(c_{j})$ or $c_{j}$ prefers $s_{i}$ over some $s_{k}$ assigned to $c_{j}$. That is, for every $c_{j}\in B$, either $|S_{M}(c_{j})|<q^{+}(c_{j})$ or $s_{i}\succ_{c_{j}}s_{k}$ for some $s_{k}\in S_{M}(c_{j})$.
  2. 2.
    There exists a subset $D_{i}\subseteq C_{M}(s_{i})$ (possibly empty) such that $s_{i}$ prefers her most preferred course in $B$ to her most preferred course in $D_{i}$, $O(B)\geq O(D_{i})$ and $O(C_{M}(s_{i}))-O(D_{i})+O(B)\leq T(s_{i})$. That is, $s_{i}$ has enough credits remaining for all courses in $B$, possibly after removing a subset of courses $D_{i}$ with a total number of credits smaller than or equal to that of $B$ such that $s_{i}$ prefers her most preferred course in $B$ over her most preferred course in $D_{i}$.
  3. 3.
    If there are downward-feasible constraints, the matching $(M\cup\{(s_{i},c_{j})|c_{j}\in B\})\setminus\{(s_{i},c_{k}):c_{k}\in D_{i}\}$ is feasible. That is, removing $D_{i}$ and adding $c_{j}$ to $C_{M}(s_{i})$ for every $c_{j}\in B$ does not result in an infeasible matching.
If $M$ has no first-blocking coalition then it is first-coalition-stable. Course Allocation with first-coalition stability is denoted ca-fc.
The four stability definitions can be ordered in terms of their strength with the aid of the following proposition.
Proposition 1.
Pair stability implies first-coalition stability, which implies coalition stability, which implies pair-size stability.
Proof.
We will prove all three implications by contradiction. First, assume that a matching $M$ is pair-stable but not first-coalition-stable. Hence, there exist a student $s_{i}$ and a set of courses $B$ such that $s_{i}$ has enough space for all courses in $B$, possibly after removing a subset of courses $D_{i}\subseteq C_{M}(s_{i})$ that $s_{i}$ prefers less than her most preferred course in $B$. Let this course be $c_{j}$. Then $(s_{i},c_{j})$ is a blocking pair, so $M$ is not pair-stable, a contradiction. If there are downward-feasible constraints, $(s_{i},c_{j})$ is still a blocking pair, because $s_{i}$ and $B$ form a first-blocking coalition, which means that $(M\cup\{(s_{i},c_{j})|c_{j}\in B\})\setminus\{(s_{i},c_{k}):c_{k}\in D_{i}\}$ is feasible, and $(M\cup(s_{i},c_{j}))\setminus\{(s_{i},c_{k}):c_{k}\in D_{i}\}$ is a subset of that matching, so also feasible.
Now, assume that $M$ is first-coalition-stable, but not coalition-stable. Then there exist a student $s_{i}$ and a set of courses $B$ such that $s_{i}$ has enough space for all courses in $B$, possibly after removing a subset of courses $D_{i}\subseteq C_{M}(s_{i})$ that $s_{i}$ prefers less than all courses in $B$. This implies that $s_{i}$ prefers her most preferred course in $B$ over her most preferred course in $D_{i}$, so $s_{i}$ and $B$ form a first-blocking coalition, and $M$ is not first-coalition-stable, a contradiction. If there are downward-feasible constraints, $s_{i}$ and $B$ is still a first-blocking coalition, because $s_{i}$ and $B$ form a blocking coalition, which means that $(M\cup\{(s_{i},c_{j})|c_{j}\in B\})\setminus\{(s_{i},c_{k}):c_{k}\in D_{i}\}$ is feasible.
Finally. assume that $M$ is coalition-stable, but not pair-size-stable. Then there exist a student $s_{i}$ and a course $c_{j}$ such that $(s_{i},c_{j})$ is a size-blocking pair. However, this implies that $s_{i}$ and $\{c_{j}\}$ form a blocking coalition, so $M$ is not coalition-stable, a contradiction. If there are downward-feasible constraints, $s_{i}$ and $\{c_{j}\}$ is still a blocking coalition, because $(s_{i},c_{j})$ is a size-blocking pair, which means that $(M\cup(s_{i},c_{j}))\setminus\{(s_{i},c_{k}):c_{k}\in D_{i}\}$ is feasible. ∎
3 Verifying stability
In this section we consider the complexity of the problem of determining whether a given matching is stable. Firstly, we remark that it is fairly trivial to check whether an assignment is a matching, and whether a matching is feasible, in the presence of downward-feasible constraints. One simply needs to check if the assignment respects upper quotas and credit limits, and if it respects all downward-feasible constraints. On the other hand, checking whether a (feasible) matching is stable may be trickier, depending on the stability definition. Let ca-x-test be the problem of deciding whether a given matching $M$ is stable, for a given definition of stability x. Under pair stability, one can verify the stability of $M$ in polynomial time. For any acceptable student-course pair $(s_{i},c_{j})\notin M$, we can check if $c_{j}$ has enough capacity for $s_{i}$, or if $c_{j}$ prefers $s_{i}$ over some $s_{k}\in S_{M}(c_{j})$, which takes $O(n)$ time, where $n$ is the number of agents. Similarly, we can check in $O(m)$, where $m$ is the number of courses, if $s_{i}$ has enough remaining credits for $c_{j}$, possibly after removing all courses in $C_{M}(s_{i})$ that $s_{i}$ prefers less than $c_{j}$. Recall that under this definition a student can end up with fewer credits after satisfying a blocking pair. As there are $O(nm)$ acceptable pairs, the total complexity is $O(n^{2}m^{2})$. Thus, we obtain the following observation.
Observation 2.
ca-p-test is solvable in $O(n^{2}m^{2})$ time, where $n=|S|$ and $m=|C|$.
However, verifying stability is hard for the other three definitions of stability, which we will show in Theorem 1. The difference is that in a blocking pair (in pair stability), a student is allowed to end up with fewer credits than she had in the matching $M$, but that is not allowed in the other three definitions of stability. This fact is crucial in our transformation from the Subset Sum problem. An instance of Subset Sum consists of a multiset of integers $X$ and a target integer $T$. The goal is to determine the existence of a subset of $X$ where the sum of its elements equals $T$. The problem is weakly NP-complete, even if every element in $X$ is a positive integer (garey_1979_computers, , page 223).
Theorem 1.
The following problems are weakly co-NP-complete: ca-fc-test, ca-c-test and ca-ps-test.
Proof.
These problems are in co-NP because given a matching $M$, a student $s$, a set of courses $B$ (of size $1$ in the case of pair-size stability) with which $s$ forms a size-blocking pair/blocking coalition/first-blocking coalition, and a set of courses $D$ that $s$ would drop, we can check whether $s$ blocks the matching $M$ with the courses in $B$.
Consider a Subset Sum instance $I$, with multiset $X$ and target number $T$. We create an instance $J$ of Course Allocation as follows. For every element $e\in X$, we create a course $c_{e}$, with upper quota $1$ and $O(c_{e})=e$. We create one additional course, $b$, with upper quota $1$ and $O(b)=T$. We also create one student, $s$, with $T(s)$ equal to the sum of all elements in $X$. The preference list of $s$ is $b$ first, then all of the other courses in any order. The matching that we want to verify is $M=\{(s,c_{e})|e\in X\}$. Thus, the matching includes every acceptable pair except for $(s,b)$. Also, $O(C_{M}(s)=T(s))$.
Assume that there exists a subset $X^{\prime}\subseteq X$ such that $\sum_{e\in X^{\prime}}e=T$. Let $D=\{c_{e}|e\in X^{\prime}\}$ be a subset of courses. Then $(s,b)$ is a size-blocking pair, because $b$ has enough capacity for $s$, $s$ prefers $b$ over every course in $D$, $O(b)\geq O(D)$, and $O(C_{M}(s))-O(D)+O(b)\leq T(s)$. Thus, $M$ is not pair-size-stable. We have that $s$ and $\{B\}$ form a blocking coalition and a first-blocking coalition, too, so $M$ is not coalition-stable or first-coalition-stable either.
Now, assume that $M$ is not pair-size-stable. Thus, there exists at least one size-blocking pair. The only acceptable pair that is not in $M$ is $(s,b)$, so that is our size-blocking pair. As $(s,b)$ is a size-blocking pair, $s$ prefers $b$ over a subset of courses $D\subseteq C_{M}(s)$ such that $O(b)\geq O(D)$ and $O(C_{M}(s))-O(D)+O(b)\leq T(s)$. As $T(s)-O(C_{M}(s))+O(D)=O(b)$, we have that $O(b)=T=O(D)$. As $b\notin D$, all courses in $D$ are of type $c_{e}$. Thus, we can build a subset $X^{\prime}\subseteq X$, such that $X=\{e|c_{e}\in D\}$ and $\sum_{e\in X^{\prime}}e=T$. Therefore, there exists a subset of $X$ whose elements add up to $T$. This direction also works in the case of coalition stability or first-coalition stability. ∎
However, just like with Subset Sum, a pseudo-polynomial dynamic programming algorithm can verify the stability of a given matching, for either pair-size stability, coalition stability, or first-coalition stability. Thus, if the number of credits of all courses is not exponentially large, verifying the stability of a matching can be done in polynomial time. In real-life settings, it is reasonable to assume that course credits are polynomially bounded.
As an aside, verifying the stability of a matching is weakly co-NP-complete even in an instance of ca-ml, where students are ordered in a master list. In the proof of Theorem 1 there is a single student, who trivially forms a master list on her own.
4 Finding a stable matching
4.1 Arbitrary course preferences
We now move onto the problem of finding a stable matching or reporting that none exists, for our four stability definitions. Let ca-x-find be the problem of finding a stable matching in a ca instance $I$, for a given definition of stability x, or determining that none exists. As in the case of testing, the complexity of ca-x-find depends on the definition of stability. In fact, depending on the definition of stability, a stable matching may or may not be guaranteed to exist.
Let us first consider the problem of finding a pair-size-stable matching. In order to do so, we use the following procedure. First, we group courses by number of credits, from largest to smallest. Let $\{O(c_{j}):c_{j}\in C\}=\{N_{1},\dots,N_{R}\}$ for some $R\geq 1$, where $N_{k}>N_{k+1}$ ($11\leq k\leq R-$ 1). Let $C_{k}=\{c_{j}\in C:O(c_{j})=N_{k}\}$.
A pair-size-stable matching can be found using a variation of Deferred Acceptance, which has $R$ rounds. In the first round, we consider the set of all students $S$ and the subset $C_{1}$ of courses, and match them using DA. We take this partial matching onto round $2$, in which we consider all students $S$ and the subset $C_{2}$ of courses. In general, in round $i$, when we consider all students $S$ and the subset $C_{i}$ of courses, each student applies to courses in $C_{i}$ from the beginning of their preference list. Importantly, if a student $s_{i}$ and a course $c_{j}$ are matched in some round, they remain matched until the end of execution, so at each round the number of credits that $s_{i}$ has available for courses stays the same or decreases. We continue until $R$ rounds have been executed, at which point the entire set $C=\bigcup_{i=1}^{R}C_{i}$ of courses will have been considered. The pseudocode for this algorithm is given in Algorithm 1.
Algorithm 1 Algorithm for finding a pair-size-stable (feasible) matching
0:  Course Allocation instance $I$
0:  return a pair-size-stable matching $M$ in $I$
1:  $M:=\emptyset$;
2:  for $k\in 1\dots R$ do
3:     while some student $s_{i}\in S$ satisfies $O(C_{M}(s_{i}))+N_{k}\leq T(s_{i})$ and $s_{i}$ has not applied to all feasible courses in $C_{k}$ on her list do
4:        $c_{j}:=$ $s_{i}$’s most preferred feasible course in $C_{k}$ that $s_{i}$ has not yet applied to;
5:        $s_{i}$ applies to $c_{j}$;
6:        if $|S_{M}(c_{j})|<q^{+}(c_{j})$ then
7:           $M:=M\cup\{(s_{i},c_{j})\}$;
8:        else
9:           $s_{k}:=$ worst student assigned to $c_{j}$;
10:           if $c_{j}$ prefers $s_{i}$ to $s_{k}$ then
11:              $M:=(M\cup\{(s_{i},c_{j})\})\backslash\{(s_{k},c_{j})\}$;
12:              $c_{j}$ rejects $s_{k}$;
13:           else
14:              $c_{j}$ rejects $s_{i}$;
15:           end if
16:        end if
17:     end while
18:  end for
19:  return  $M$;
It is clear that Algorithm 1 finds a matching $M$. Furthermore, if there are downward-feasible constraints, $M$ is feasible, because every $s_{i}$ applies only to courses $c_{j}$ such that adding $(s_{i},c_{j})$ to $M$ does not violate downward-feasible constraints. With appropriate data structures, the total complexity of Algorithm 1 is $O(nm^{2})$, where $n=|S|$ and $m=|C|$. This is because for every student we consider each course in her preference list at most $m$ times, if there are downward-feasible constraints, as courses that were previously unfeasible might become feasible for some $s_{i}$ if she is rejected from some $c_{j}$. Meanwhile, when a student applies to a course, the course can perform all necessary steps in constant time, giving a total time complexity of $O(nm^{2})$, where $n=|S|$ and $m=|C|$. If there are no downward-feasible constraints then the time complexity is $O(nm)$. We now show that $M$ is pair-size-stable.
Theorem 2.
Algorithm 1 always finds a pair-size-stable (feasible) matching in $O(nm^{2})$ time, where $n=|S|$ and $m=|C|$.
Proof.
We will argue by contradiction. Assume the matching $M$ found by Algorithm 1 is not pair-size-stable, and that there is a size-blocking pair $(s_{i},c_{j})$, such that $c_{j}\in C_{l}$ for some $l$. Thus, either $s_{i}$ has enough remaining credits to match with $c_{j}$ (and adding $(s_{i},c_{j})$ to $M$ results in a feasible matching) or $s_{i}$ prefers $c_{j}$ over every course $c_{k}$ from a subset of courses $D_{i}$ such that $s_{i}$ is assigned to every $c_{k}\in D_{i}$, $O(c_{j})\geq\sum_{c_{k}\in D_{i}}O(c_{k})$, and, if there are downward-feasible constraints, $(M\cup\{(s_{i},c_{j})\})\backslash D_{i}$ is feasible.
Assume that the first case holds. The course $c_{j}$ is in subset $C_{l}$, meaning that it was considered in the $l$-th round. As $s_{i}$ has enough credits for $c_{j}$ after the execution of the algorithm, it had enough credits during round $l$, because matches between students and courses cannot be broken, so after every round the number of remaining credits for $s_{i}$ stays the same or decreases. Thus, $s_{i}$ applied to $c_{j}$ during round $l$, so if they are not matched at the end of round $l$ then $c_{j}$ rejected $s_{i}$ for a better student. Hence, $(s_{i},c_{j})$ was not a size-blocking pair at the end of round $l$. Because all matches formed in round $l$ are part of the final pair-size-stable matching, $c_{j}$ prefers all of its students over $s_{i}$, so $(s_{i},c_{j})$ is not a size-blocking pair, a contradiction. Therefore, this case cannot occur.
Now, assume the second case holds, so $s_{i}$ prefers $c_{j}$ over every course $c_{k}$ from a subset of courses $D_{i}$ such that $s_{i}$ is assigned to every $c_{k}\in D_{i}$ and $O(c_{j})\geq\sum_{c_{k}\in D_{i}}O(c_{k})$. As $O(c_{j})\geq\sum_{c_{k}\in D_{i}}O(c_{k})$, $O(c_{j})\geq O(c_{k})$ $\forall c_{k}\in D_{i}$, so no $c_{k}\in D_{i}$ has a number of credits strictly greater than $O(c_{j})$. Additionally, no $c_{k}\in D_{i}$ has a number of credits equal to $O(c_{j})$, for if there were such a $c_{k}$ and $(s_{i},c_{j})$ was a size-blocking pair, then that would imply that in the round where $c_{j}$ and $c_{k}$ were considered, Deferred Acceptance did not find a stable matching, which is a contradiction. Therefore, $c_{j}$ has a number of credits strictly greater than every $c_{k}\in D_{i}$. However, this means that, during the algorithm’s execution, $c_{j}$ was considered in an earlier round than every $c_{k}\in D_{i}$, round $l$. During that round, $s_{i}$ had enough remaining credits for $c_{j}$, because, after removing every course from $D_{i}$, it has enough credits for $c_{j}$ after the execution of the algorithm. Using a similar argument to the first case, $(s_{i},c_{j})$ was not a size-blocking pair at the end of round $l$, meaning that it is not a size-blocking pair upon termination of the algorithm, a contradiction. Therefore, this case also cannot occur. By assuming either case holds, we arrive at a contradiction, which is caused by the assumption that there exists a size-blocking pair. Therefore, no such pair exists, and the matching is pair-size-stable. ∎
On the other hand, a first-coalition-stable matching, or pair-stable matching, may not always exist. Consider, for example, the instance in Example 1.
Example 1.
Let us look at every possible matching in the instance from Figure 4. Then for every matching $M$ we can find a first-blocking coalition, which implies that $M$ is not first-coalition-stable, and thus that it is not pair-stable.
  • If $M=\{(s_{1},c_{1}),(s_{1},c_{2})\}$, then $(s_{2},c_{1})$ is a first-blocking coalition.
  • If $M=\{(s_{1},c_{1}),(s_{2},c_{2})\}$, then $(s_{1},c_{2})$ is a first-blocking coalition.
  • If $M=\{(s_{1},c_{2}),(s_{2},c_{1})\}$, then $(s_{1},c_{3})$ is a first-blocking coalition.
  • If $M=\{(s_{1},c_{3}),(s_{2},c_{2})\}$, then $(s_{1},\{c_{1},c_{2}\})$ is a first-blocking coalition.
  • If $M=\{(s_{1},c_{3}),(s_{2},c_{1})\}$, then $(s_{2},c_{2})$ is a first-blocking coalition.
$s_{i}$ $T(s_{i})$ $\succ_{s_{i}}$ $c_{j}$ $O(c_{j})$ $q^{+}(c_{j})$ $\succ_{c_{j}}$
$s_{1}$ 2 $c_{1}$ $\succ$ $c_{3}$ $\succ$ $c_{2}$ $c_{1}$ 1 1 $s_{2}$ $\succ$ $s_{1}$
$s_{2}$ 1 $c_{2}$ $\succ$ $c_{1}$ $c_{2}$ 1 1 $s_{1}$ $\succ$ $s_{2}$
$c_{3}$ 2 1 $s_{1}$
Figure 4: Instance of Course Allocation. This instance is derived from a Hospitals/Residents with Sizes instance due to McDermid and Manlove mcdermid_keeping_2010 .
Determining whether a pair-stable matching exists is NP-complete, which can be seen by considering a connection between course allocation and the problem Hospitals/Residents with Sizes, a modification of Hospitals/Residents in which some residents need more than one space mcdermid_keeping_2010 . With respect to the stability definition given by McDermid and Manlove mcdermid_keeping_2010 , the problem of deciding whether a stable matching exists, given an instance of Hospitals/Residents with Sizes, is NP-complete. Moreover this result holds even if each resident has size 1 or 2, and each hospital has upper quota at most 2. By replacing residents by courses and hospitals by students, we can then obtain a simple reduction to course allocation where each student $s_{i}$ satisfies $T(s_{i})\leq 2$ and each course $c_{j}$ satisfies $O(c_{j})\in\{1,2\}$. Moreover, McDermid and Manlove’s stability definition corresponds to pair stability. The following result is therefore immediate.
Theorem 3 (mcdermid_keeping_2010 ).
ca-p-find is NP-hard, even if there are no downward-feasible constraints.
Similarly, determining the existence of a first-coalition-stable matching is NP-hard, and we show that by transforming from Hospitals/Residents with Sizes, though in this case a more complex reduction is required.
Theorem 4.
ca-fc-find is NP-hard, even if there are no downward-feasible constraints.
Proof.
We prove this result by transforming from the problem of determining the existence of a stable matching in the Hospitals/Residents with Sizes (HRS) setting, which was shown to be NP-complete by McDermid and Manlove mcdermid_keeping_2010 . In the HRS instance $I$ that we are transforming from, each resident has size $1$ or $2$, each hospital has upper quota $2$, and the length of every preference list is at most $3$.
The transformation proceeds as follows. Every resident $r_{j}$ in $I$ becomes a course $c_{j}$ in the Course Allocation instance $J$, with upper quota $1$, and number of credits $1$ or $2$, depending on the size of $r_{j}$. Every hospital $h_{i}$ becomes a student $s_{i}$, with $T(s_{i})=2$. We use this bijection to transform every preference list. For example, if resident $r_{j}$ has preference list $h_{i}\succ_{r_{j}}h_{k}\succ_{r_{j}}h_{l}$, now course $c_{j}$ has preference list $s_{i}\succ_{c_{j}}s_{k}\succ_{c_{j}}s_{l}$. Additionally, for every student $s_{i}$ we create a dummy course $c_{i}^{\prime}$, of size $1$, and append this course to the end of $s_{i}$’s preference list. The preference list of $c_{i}^{\prime}$ contains $s_{i}$ only.
Assume that the HRS instance $I$ has a stable matching $M$. Then we build a first-coalition-stable matching $M^{\prime}$ in $J$ as follows. We take $M$ and swap residents and hospitals for courses and students, respectively. Then, for every student $s_{i}$ such that $C_{M^{\prime}}(s_{i})<2$, we add the pair $(s_{i},c_{i}^{\prime})$ to $M$. This matching is first-coalition-stable. Assume it is not, which implies the existence of a first-blocking coalition involving one student $s_{i}$. Let us go through each possibility:
  • Student $s_{i}$ has only $1$ credit (which must be from course $c_{i}^{\prime}$, according to the construction of $M^{\prime}$), and forms a first-blocking coalition with a course $c_{j}$, with size $1$ or $2$. In this case, hospital $h_{i}$ would have no residents in $M$, and resident $r_{j}$ would prefer $h_{i}$ over her other partner, so $(r_{j},h_{i})$ would block $M$ in $I$, a contradiction.
  • Student $s_{i}$ has $2$ credits, one of which comes from $c_{i}^{\prime}$ and the other from some $c_{k}$, and she forms a first-blocking coalition with course $c_{j}$. If $c_{j}$ has size $1$, then $s_{i}$ prefers $c_{j}$ over $c_{i}^{\prime}$. However, this implies that in $M$ hospital $h_{i}$ had one resident only, with capacity for resident $r_{j}$, which preferred $h_{i}$ over her assigned hospital, forming a blocking pair, a contradiction. If $c_{j}$ has size $2$, then $s_{i}$ prefers $c_{j}$ over $c_{k}$, which implies that in $I$ hospital $h_{i}$ prefers resident $r_{j}$ over its assigned resident $r_{k}$, and $r_{j}$ prefers $h_{i}$ over her assigned hospital, forming a blocking pair, a contradiction.
  • Student $s_{i}$ has $2$ credits, which come from two courses of size $1$, namely $c_{p}$ and $c_{q}$, or one course of size $2$, namely $c_{p}$, and she forms a first-blocking coalition with either one course $c_{j}$, or with two courses, $c_{j}$ and $c_{k}$, such that $c_{j}\succ_{s_{i}}c_{k}$. Then, in $I$, hospital $h_{i}$ prefers resident $r_{j}$ over either resident $r_{p}$ or residents $r_{p}$ and $r_{q}$, and $r_{j}$ prefers $h_{i}$ over her assigned hospital. Hence, $(r_{j},h_{i})$ blocks $M$ in $I$, a contradiction.
Similarly, assume that our Course Allocation instance $J$ has a first-coalition-stable matching $M^{\prime}$. Then we take $M^{\prime}$, remove any pair containing a course of the form $c_{i}^{\prime}$, and swap courses and students for residents and hospitals to get a matching $M$ in the HRS instance $I$. This matching $M$ is also stable in $I$: if $(r_{j},h_{i})$ were to block $M$ in $I$, then either $(s_{i},c_{j})$ or $(s_{i},\{c_{j},c_{i}^{\prime}\})$ would block $M$ in $J$, a contradiction.
Therefore, there exists a first-coalition-stable matching in $J$ if and only if there exists a stable matching in $I$. Thus, as HRS is NP-complete, so is ca-fc-find. ∎
4.2 Master lists
An important restriction of Course Allocation arises when students are ordered in a master list, for example by the average of their grades. Let ca-x-ml-find be the problem of finding a stable matching in a ca-ml instance $I$, for a given definition of stability x, or determining that none exists. When students are ranked according to a master list, we can use the Serial Dictatorship mechanism to find a stable matching, for any definition of stability. Serial Dictatorship is an algorithm that starts with an empty matching $M$ and, on round $i$, considers the $i$-th student in the master list, say $s_{i}$. Then, it considers every course on $s_{i}$’s preference list, from most to least preferred. For each such $c_{j}\in P(s_{i})$, it checks whether $c_{j}$ has some capacity left, whether adding $(s_{i},c_{j})$ to $M$ violates the credit limit of $s_{i}$, and, if there are downward-feasible constraints, whether adding $(s_{i},c_{j})$ to $M$ violates any of them. If it is possible to add $(s_{i},c_{j})$ to $M$, it does so, and proceeds to check the next course in $P(s_{i})$. This is clearly a $O(nm)$ time algorithm, where $n=|S|$ and $m=|C|$. We now show that Serial Dictatorship always finds a pair-stable (feasible) matching.
Proposition 2.
In any ca-ml instance $I$, a pair-stable matching always exists, and one can use Serial Dictatorship to find such a matching in $O(nm)$ time, where $n=|S|$ and $m=|C|$.
Proof.
Aiming for a contradiction, assume that after executing Serial Dictatorship, student $s_{i}$ and course $c_{j}$ form a blocking pair. Then this means that either $c_{j}$ has some capacity left for $s_{i}$ or $c_{j}$ prefers $s_{i}$ over some $s_{k}$ that is assigned to $c_{j}$. If the second case holds, $s_{i}\succ_{ML}s_{k}$, so $s_{i}$ got to choose courses before $s_{k}$, meaning that in the round of $s_{i}$, $s_{k}$ was not assigned to $c_{j}$ and $c_{j}$ had an open place that $s_{i}$ could have taken. If $s_{i}$ did not choose $c_{j}$ and instead chose other courses, then $s_{i}$ preferred her assigned courses over $c_{j}$, which contradicts the assumption that $s_{i}$ and $c_{j}$ form a blocking pair. The reasoning is similar if the first case holds instead. Therefore, the algorithm above always produces a pair-stable matching. ∎
By Proposition 1, a pair-stable matching is first-coalition-stable, coalition-stable, and pair-size-stable. Therefore, the matching found by Serial Dictatorship is first-coalition-stable, coalition-stable, and pair-size-stable, and thus ca-x-ml-find is solvable in polynomial time for any definition of stability x.
Moreover, the pair-stable matching found by Serial Dictatorship is the unique pair-stable matching. Assume otherwise, so that there are two different pair-stable matchings, $M$ and $M^{\prime}$, and let $s_{i}$ be the highest-ranked student such that $C_{M}(s_{i})\neq C_{M^{\prime}}(s_{i})$. That is, she is assigned different courses in each matching. Let $c_{j}$ be the most preferred course of $s_{i}$ in $C_{M}(s_{i})\ominus C_{M^{\prime}}(s_{i})$, the symmetric difference of $C_{M}(s_{i})$ and $C_{M^{\prime}}(s_{i})$. Without loss of generality, assume that $c_{j}\in C_{M}(s_{i})$. Then $(s_{i},c_{j})$ blocks $M^{\prime}$, because $s_{i}$ prefers to drop all courses in $C_{M^{\prime}}(s_{i})\setminus C_{M}(s_{i})$ and be allocated to $c_{j}$ instead, and the set $(C_{M}(s_{i})\cap C_{M^{\prime}}(s_{i}))\cup\{(s_{i},c_{j})\}$ is feasible, as it is a subset of the feasible set $C_{M}(s_{i})$. Meanwhile, because $s_{i}$ is the highest-ranked student that has different courses in $M$ and $M^{\prime}$, $c_{j}$ prefers $s_{i}$ over the students that it has in $M^{\prime}$. Hence, we obtain a contradiction to the pair stability of $M^{\prime}$, which implies that there is a unique pair-stable matching.
Observation 3.
Every instance of ca-ml has a unique pair-stable matching.
Under other definitions of stability, an instance of ca-ml might have more than one stable matching. For example, consider an instance with two courses, $c_{1}$ and $c_{2}$, such that $O(c_{1})=1$ and $O(c_{2})=2$, and one student, $s_{1}$, with $c_{1}\succ_{s_{1}}c_{2}$ and $T(s_{1})=2$. Then both $\{(s_{1},c_{1})\}$ and $\{(s_{1},c_{2})\}$ are first-coalition-stable matchings, which implies that they are coalition-stable and pair-size-stable, too. $\{(s_{1},c_{2})\}$ is not a pair-stable matching though, because $s_{1}$ prefers $c_{1}$ and under pair stability she can end up with fewer credits when participating in a blocking pair.
5 Finding a maximum size stable matching
5.1 No lower quotas
As we saw in the previous section, every instance of ca-ml has at least one stable matching, for any definition of stability. However, there may be more than one possible stable matching, and indeed they can have different sizes, in which case we want to find a stable matching that is as large as possible. We refer to this problem as ca-x-ml-max, where x is a definition of stability. ca-p-ml-max is trivially solvable in polynomial time, by Observation 3.
However, under the other three definitions of stability, ca-x-ml-max is NP-hard. In fact, under coalition stability, ca-c-ml-max is NP-hard even if courses and students are ordered in two master lists, every course has $1$ or $2$ credits and upper quota $1$, and every student can take up to $2$ credits.
In order to prove this result, we will transform from the Stable Marriage problem with Ties and Incomplete preferences (smti). com-smti, the problem of determining whether a complete weakly stable matching in an smti instance exists, is NP-complete even if there are master lists on both sides, the ties occur in one master list only, and are of length $2$ (irving_stable_2008, , Thm. 3.2).
The key to understanding the transformation from Theorem 5, which shows that ca-c-ml-max is NP-hard, is in Figure 5. In this instance, $c_{2}$ is a $2$-credit course and $c_{1}$ and $c_{3}$ are $1$-credit courses. Thus, under coalition stability, $s_{1}$ is indifferent between being matched with $c_{1}$ and $c_{3}$, and with $c_{2}$. Hence, under coalition stability, we can speak of a tie between $c_{2}$ on the one hand, and $c_{1}$ and $c_{3}$ on the other, even though preference lists are all strict. This idea gives the connection between com-smti and ca-c-ml-max.
$s_{1}$ $:c_{1}\succ_{s_{1}}c_{2}\succ_{s_{1}}c_{3}$
Figure 5: Preference list of $s_{1}$, with $T(s_{1})=2$. $O(c_{1})=O(c_{3})=1$, and $O(c_{2})=2$.
Theorem 5.
It is NP-hard to determine whether a ca-ml instance has a student-complete (or course-complete) coalition-stable matching, even if both students and courses are ordered according to master lists, every student takes up to $2$ credits, and every course is of size $1$ or $2$ with upper quota $1$. Therefore, ca-c-ml-max is NP-hard, even if there are no downward-feasible constraints.
Proof.
We transform from the following restriction of SMTI: in the SMTI instance $I$, both men and women are ordered in two master lists, ties only occur in the women’s master list (so on men’s preference lists only), and each tie is of length $2$. Under these conditions, COM-SMTI is NP-complete. We construct a ca-ml instance $J$ from $I$.
Every woman $w_{j}$ becomes a course $c_{j}$, with the same preference list, and with upper quota $1$. If $w_{j}$ is not involved in a tie in the women’s master list then $c_{j}$ is a $2$-credit course. If women $w_{k},w_{l}$ are tied in the women’s master list, then one of them (say $w_{k}$) becomes a $1$-credit course ($c_{k}$), while the other ($w_{l}$) becomes a $2$-credit course ($c_{l}$). For every $1$-credit course $c_{k}$ we create another $1$-credit course $c_{k}^{\prime}$. This course has the same preference list as $c_{k}$, substituting men for students, and it also has upper quota $1$. The master list of courses is the same as the master list of women, except that if women $w_{k},w_{l}$ are tied in the women’s master list, then the master list of courses (MLC) includes $c_{k}\succ_{MLC}c_{l}\succ_{MLC}c_{k}^{\prime}$, where $O(c_{k})=O(c_{k}^{\prime})=1$ and $O(c_{l})=2$.
Every man $m_{i}$ becomes a student $s_{i}$, who can take up to $2$ credits. The preferences of $s_{i}$ are inherited from $m_{i}$, such that every woman $w_{j}$ not involved in a tie in the women’s master list becomes $c_{j}$ in $s_{i}$’s preference list. If some $w_{j}$ is in a tie with $w_{k}$ in the women’s master list, but only $w_{j}$ is in the preference list of $m_{i}$, then we replace $w_{j}$ by $c_{j}$, if $c_{j}$ is a $2$-credit course, or by $c_{j}$ first and then $c_{j}^{\prime}$, if $c_{j}$ is a $1$-credit course. If both $w_{j}$ and $w_{k}$ are tied and in $m_{i}$’s preference list, then we substitute them both by $c_{j}\succ_{s_{i}}c_{k}\succ_{s_{i}}c_{j}^{\prime}$. By Figure 5, we know that $s_{i}$ is indifferent between being matched with $\{c_{j},c_{j}^{\prime}\}$ and with $\{c_{k}\}$, under coalition stability. From the men’s master list we derive a students’ master list. We now show that course-complete matchings are student-complete, and vice versa.
Lemma 1.
Let $M^{\prime}$ be a matching in $J$. Then $M^{\prime}$ is course-complete if and only if it is student-complete.
Proof.
Let $n$ be the number of men (and women) in the SMTI instance $I$. Then, in the ca-ml instance $J$, for each woman $w_{j}$ we have either one $2$-credit course with upper quota $1$, or two $1$-credit courses with upper quota $1$. Thus, a course-complete matching has a total of $2\cdot n$ credits. Similarly, for each man we have a student $s_{i}$ with $T(s_{i})=2$. Thus, a student-complete matching has $2\cdot n$ credits. Therefore, a course-complete matching has the same number of credits as a student-complete matching. ∎
A consequence of this transformation is that, in a coalition-stable matching $M^{\prime}$, a student $s_{i}$ is matched with either a $2$-credit course $c_{j}$ or two $1$-credit courses $\{c_{k},c_{k}^{\prime}\}$, such that both courses come from the same woman $w_{k}$. Assume that this is not the case, and that there exists some $s_{i}$ that is matched with a $1$-credit course $c_{j}$ but not with $c_{j}^{\prime}$ (the other case, where $s_{i}$ is matched to $c_{j}^{\prime}$ instead, can be proved analogously). Furthermore, assume that $s_{i}$ is the highest placed student in this situation and that $c_{j}$ is ranked higher than the other course matched to $s_{i}$, if there is one. Then, because $s_{i}$ is the highest ranked student matched to one but not both of the courses that come from the same woman, $c_{j}^{\prime}$ prefers $s_{i}$ over its only student. Similarly, if $s_{i}$ is matched to some $1$-credit course $c_{l}$, then she prefers $c_{j}^{\prime}$ over $c_{l}$. Therefore, $(s_{i},\{c_{j}^{\prime}\})$ is a blocking coalition, a contradiction to the stability of $M^{\prime}$. Hence, if a student $s_{i}$ is matched with a $1$-credit course $c_{j}$, she is also matched with $c_{j}^{\prime}$.
Now we show that there exists a complete weakly stable matching in $I$ if and only if there exists a student-complete (course-complete) coalition-stable matching in $J$. First, assume that there is a complete weakly stable matching $M$ in $I$. Then we build a complete matching $M^{\prime}$ in $J$ as follows. For each $(m_{i},w_{j})\in M$ such that $c_{j}$ is a $2$-credit course, add $(s_{i},c_{j})$ to $M^{\prime}$. For each $(m_{i},w_{j})\in M$ such that $c_{j}$ is a $1$-credit course, add $(s_{i},c_{j})$ and $(s_{i},c_{j}^{\prime})$ to $M^{\prime}$. Clearly the matching is both student-complete and course-complete, as every course is matched, and every student takes $2$ credits. We now show that $M^{\prime}$ is coalition-stable. Assume that it is not, and that there exists a blocking coalition $(s_{p},\{c_{q}\})$, with $O(c_{q})=2$ (or $(s_{p},\{c_{q},c_{q}^{\prime}\})$ where $O(c_{q})=1$). Thus, $s_{p}$ strictly prefers $c_{q}$ (and $c_{q}^{\prime}$, if it exists) over every course in $C_{M^{\prime}}(s_{p})$. This implies that $c_{q}$ (and $c_{q}^{\prime}$, if it exists) is placed higher in the courses’ master list than every course in $C_{M^{\prime}}(s_{p})$. Similarly, $c_{q}$ (and $c_{q}^{\prime}$, if it exists) strictly prefers $s_{p}$ over its student $s_{r}$, meaning that $s_{p}$ is placed higher in the students’ master list than $s_{r}$. However, this means than in $I$, $w_{q}$ prefers $m_{p}$ over $m_{r}$, and that $m_{p}$ prefers $w_{q}$ over his partner in $M$, meaning that $(m_{p},w_{q})$ blocks $M$, a contradiction to the stability of $M$. Thus, $M^{\prime}$ is a coalition-stable matching.
Now, assume that $J$ has a student-complete (course-complete) coalition-stable matching $M^{\prime}$. Thus, every student $s_{i}$ is matched with either a $2$-credit course $c_{j}$, or two $1$-credit courses, $c_{k}$ and $c_{k}^{\prime}$. We build a complete matching $M$ in $I$ as follows. If $(s_{i},c_{j})\in M^{\prime}$, add $(m_{i},w_{j})$ to $M$. Clearly $M$ is complete, because every man and every woman is matched. We now show that $M$ is weakly stable. Assume that it is not, so that there exists a blocking pair $(m_{p},w_{q})$. Thus, $m_{p}$ prefers $w_{q}$ over his partner in $M$, and $w_{q}$ prefers $m_{p}$ over her partner, which means that $w_{q}$ and $m_{p}$ are ranked higher in their respective master lists than their partners. This implies that $c_{q}$ (or $c_{q}$ and $c_{q}^{\prime}$) and $s_{p}$ are ranked higher in their respective master lists than the courses in $C_{M^{\prime}}(s_{p})$ and the only student in $S_{M^{\prime}}(c_{q})$, respectively. Hence, $s_{p}$ strictly prefers $c_{q}$ (and $c_{q}^{\prime}$, if it exists) over every course in $C_{M^{\prime}}(s_{p})$. As $c_{q}$ (and $c_{q}^{\prime}$) also strictly prefers $s_{p}$ over its only assigned student, $(s_{p},\{c_{q}\})$ (or $(s_{p},\{c_{q},c_{q}^{\prime}\})$) is a blocking coalition, which contradicts the stability of $M^{\prime}$. Thus, $M$ is a weakly stable matching.
Therefore, there exists a complete weakly stable matching in $I$ if and only if there exists a student-complete (course-complete) coalition-stable matching in $J$. Hence, determining whether a ca-ml instance has a student-complete (course-complete) coalition-stable matching is NP-hard. Thus, ca-c-ml-max is NP-hard. ∎
Likewise, determining whether a ca-ml instance has a course-complete first-coalition-stable matching is NP-hard, as well as determining whether it has a course-complete pair-size-stable matching. Hence, both ca-fc-ml-max and ca-ps-ml-max are NP-hard. However, in the following proof, we assume that any two courses have a different number of credits.
Theorem 6.
It is NP-hard to determine whether there exist a course-complete first-coalition-stable matching, even if both students and courses are ordered according to master lists, and every course has upper quota $1$. Hence, both ca-fc-ml-max and ca-ps-ml-max are NP-hard, even if there are no downward-feasible constraints.
Proof.
We reduce from COM-SMTI. Recall that COM-SMTI is NP-complete even if there are master lists on both sides, the ties occur in one master list only (in this case the women’s list), and are of length $2$. The ca-ml instance is $J$.
Every woman $w_{j}$ becomes a course $c_{j}$, with the same preference list, and with upper quota $1$. The first course in the master list, $c_{1}$, has number of credits $O(c_{1})=1+n\cdot\varepsilon$, where $n$ is the number of courses and $\varepsilon$ is an arbitrarily small positive number. The second course, $c_{2}$, has $O(c_{2})=1+(n-1)\cdot\varepsilon$, and so on, so that the $i$-th course, $c_{i}$, has number of credits $O(c_{i})=1+(n-i+1)\cdot\varepsilon$. If two courses, $c_{j}$ and $c_{j+1}$, are tied, one (say $c_{j}$) has number of credits $O(c_{j})=1+(n-j+1)\cdot\varepsilon$ and the other, $c_{j+1}$, has number of credits $O(c_{j+1})=1+(n-j)\cdot\varepsilon$. A property of this construction is that courses that are higher in the master list have a larger number of credits than those placed below. The master list of courses (MLC) mostly comes from the women’s master list. However, if two women $w_{j}$ and $w_{j+1}$ are tied in the master list, then the course with the smallest number of credits (say $c_{j+1}$) is ranked before the course with the larger number of credits (say $c_{j}$). That is, if $O(c_{j+1})<O(c_{j})$, then $c_{j+1}\succ_{MLC}c_{j}$. This way, a student matched with $c_{j}$ cannot form a first-blocking coalition with $c_{j+1}$, and vice versa.
Every man $m_{i}$ becomes a student $s_{i}$, with $T(s_{i})=1+n\cdot\varepsilon$. The preferences of $s_{i}$ are inherited from $m_{i}$, so that if $w_{j}$ is in $m_{i}$’s preference list, $c_{j}$ is in $s_{i}$’s preference list. The order of the courses in given by the master list of courses. It is easy to see that no student has enough credits for more than one course.
First, assume that there is a complete weakly stable matching $M$ in $I$. Then we build a course-complete matching $M^{\prime}$ in $J$ as follows. If $(m_{i},w_{j})\in M$, add $(s_{i},c_{j})$ to $M^{\prime}$. The matching is course-complete, as every course is matched with one student. We now show that $M^{\prime}$ is first-coalition-stable. Assume that it is not, and that there exists a first-blocking coalition $(s_{p},\{c_{q}\})$. Thus, $s_{p}$ strictly prefers $c_{q}$ over the course in $C_{M^{\prime}}(s_{p})$, and $O(c_{q})>O(C_{M^{\prime}}(s_{p}))$. This implies that $c_{q}$ is placed higher in the courses’ master list than the course in $C_{M^{\prime}}(s_{p})$. Similarly, $c_{q}$ strictly prefers $s_{p}$ over its only matched student $s_{r}$, meaning that $s_{p}$ is placed higher in the students’ master list than $s_{r}$. However, this means than in $I$ $w_{q}$ prefers $m_{p}$ over $m_{r}$, and that $m_{p}$ prefers $w_{q}$ over his partner, meaning that $(m_{p},w_{q})$ blocks $M$, a contradiction to the stability of $M$. Thus, $M^{\prime}$ is a first-coalition-stable matching.
Now, assume that $J$ has a course-complete first-coalition-stable matching $M^{\prime}$. Thus, every student $s_{i}$ is matched with one course, because no student has enough credits to be matched with two courses. We build a complete matching $M$ in $I$ as follows. For each $(s_{i},c_{j})\in M^{\prime}$, add $(m_{i},w_{j})$ to $M$. The matching is complete, as every man is matched with one woman. We now show that $M$ is stable, too. Assume that it is not, so that there exists a blocking pair $(m_{p},w_{q})$. Thus, $m_{p}$ prefers $w_{q}$ over his partner, and $w_{q}$ prefers $m_{p}$ over her partner, which means that $w_{q}$ and $m_{p}$ are ranked higher in their respective master lists than their respective partners. This implies that $c_{q}$ and $s_{p}$ are ranked higher in their respective master lists than the course in $C_{M^{\prime}}(s_{p})$ and the student in $S_{M^{\prime}}(c_{q})$, respectively. Hence, $s_{p}$ strictly prefers $c_{q}$ over the course in $C_{M^{\prime}}(s_{p})$, and $O(c_{q})>O(C_{M^{\prime}}(s_{p}))$. As $c_{q}$ also strictly prefers $s_{p}$ over its only assigned student, $(s_{p},\{c_{q}\})$ is a first-blocking coalition, which contradicts the stability of $M^{\prime}$. Thus, $M$ is a weakly stable matching.
Therefore, there exists a complete weakly stable matching in $I$ if and only if there exists a course-complete first-coalition-stable matching in $J$. Hence, determining whether a ca-ml instance has a course-complete first-coalition-stable matching is NP-hard. This same proof can be used to show that determining the existence of a course-complete pair-size-stable matching is NP-hard. ∎
In fact, the hardness of finding a maximum size first-coalition-stable matching remains even if every course has $1$ or $2$ credits. In order to prove this result, we will reduce from Min-MM. In this problem, we have a graph $G$ and a positive integer $K$, and the goal is to determine whether there exists a maximal matching of size at most $K$. This problem is NP-complete, even if $G$ is a subdivision graph horton_1993_minimum , which implies that it is bipartite and all vertices in one of the sides have degree $2$.
Theorem 7.
It is NP-hard to determine whether an instance of ca-ml has a course-complete first-coalition-stable matching, even if every student takes up to $2$ credits, every course has $1$ or $2$ credits, and each student has up to $5$ courses in her preference list.
Proof.
We reduce from Min-MM, which is NP-complete, even if $G$ is a subdivision graph horton_1993_minimum . Let $G=(U,W,E)$ be a graph representing an instance of Min-MM, where $U$ and $W$ are the two sets of vertices, $|U|=n_{1}$ and $|W|=n_{2}$. Every vertex in $W$ has degree $2$. Let $K$ be the maximum size of a maximal matching. We build an instance $I$ of ca-ml, with students $S$ and courses $C$, as follows.
For every $w_{i}\in W$, we create the following students: $s_{i}^{1}$, $s_{i}^{2}$, and $s_{i}^{3}$, with $T(s_{i}^{1})=T(s_{i}^{2})=T(s_{i}^{3})=2$, as well as $s_{i}^{4}$, with $T(s_{i}^{4})=1$. We also create courses $d_{i}^{1}$, $d_{i}^{2}$, and $d_{i}^{3}$, with $O(d_{i}^{1})=O(d_{i}^{2})=2$, and $O(d_{i}^{3})=1$. Students $s_{i}^{1}$, $s_{i}^{2}$, $s_{i}^{3}$ and $s_{i}^{4}$ belong to subset $S_{i}\subseteq S$, and courses $d_{i}^{1}$, $d_{i}^{2}$ and $d_{i}^{3}$ belong to subset $D_{i}\subseteq C$. For every $w_{i}\in W$, $s_{i}^{1}\in S^{1}\subseteq S$, $s_{i}^{2}\in S^{2}\subseteq S$, $s_{i}^{3}\in S^{3}\subseteq S$, and $s_{i}^{4}\in S^{4}\subseteq S$. Similarly, $d_{i}^{1}\in D^{1}\subseteq C$, $d_{i}^{2}\in D^{2}\subseteq C$, and $d_{i}^{3}\in D^{3}\subseteq C$.
For every $u_{j}\in U$, we create a course $c_{j}$, with $O(c_{j})=1$, and a student $p_{j}$, with $T(p_{j})=1$. Additionally, we create a course $e$, $O(e)=2$. The upper quota of every course is $1$, except for $e$, which has upper quota $n_{2}-K$. For every $u_{j}\in U$, $c_{j}\in C^{\prime}\subseteq C$ and $p_{j}\in P\subseteq S$.
The master list of students (ML) is
$s_{1}^{1}\succ_{ML}s_{1}^{2}\succ_{ML}s_{2}^{1}\succ_{ML}s_{2}^{2}\succ_{ML}...\succ_{ML}s_{n_{2}}^{1}\succ_{ML}s_{n_{2}}^{2}\succ_{ML}\\ \succ_{ML}s_{1}^{3}\succ_{ML}s_{2}^{3}\succ_{ML}...\succ_{ML}s_{n_{2}}^{3}\succ_{ML}p_{1}\succ_{ML}p_{2}\succ_{ML}...\succ_{ML}p_{n_{1}}\succ_{ML}\\ \succ_{ML}s_{1}^{4}\succ_{ML}s_{2}^{4}\succ_{ML}...\succ_{ML}s_{n_{2}}^{4}$ (1)
That is, the first students belong to subsets $S^{1}$ and $S^{2}$, such that every $s_{i}^{1}$ is followed by $s_{i}^{2}$. Then we place students from $S^{3}$, then students from $P$, and finally students from $S^{4}$, in all cases in increasing subscript order.
Let $w_{i}\in W$ be a vertex from $G$, adjacent to $u_{j}$ and $u_{k}$ (with $j<k$). We build the preference lists of $s_{i}^{1}$, $s_{i}^{2}$, $s_{i}^{3}$ and $s_{i}^{4}$ as follows.
$s_{i}^{1}$ $:c_{j}\succ_{s_{i}^{1}}d_{i}^{1}$
$s_{i}^{2}$ $:c_{k}\succ_{s_{i}^{2}}d_{i}^{1}\succ_{s_{i}^{2}}d_{i}^{2}$
$s_{i}^{3}$ $:d_{i}^{2}\succ_{s_{i}^{3}}d_{i}^{3}\succ_{s_{i}^{3}}c_{j}\succ_{s_{i}^{3}}c_{k}\succ_{s_{i}^{3}}e$
$s_{i}^{4}$ $:d_{i}^{3}$
If in the Min-MM instance $w_{i}$ is matched with $u_{j}$, then $s_{i}^{1}$ is to be matched with $c_{j}$, $s_{i}^{2}$ with $d_{i}^{1}$, $s_{i}^{3}$ with $d_{i}^{2}$ and $s_{i}^{4}$ with $d_{i}^{3}$. If, however, $w_{i}$ is unmatched, $s_{i}^{1}$ is to be matched with $d_{i}^{1}$, $s_{i}^{2}$ with $d_{i}^{2}$, $s_{i}^{3}$ with $e$ and $s_{i}^{4}$ with $d_{i}^{3}$. Student $s_{i}^{3}$ contains courses $c_{j}$, $c_{k}$, and $d_{i}^{3}$ in her preference list to make sure that a course-complete first-coalition-stable matching in $I$ corresponds to a maximal matching in $G$, so that if $s_{i}^{3}$ is matched with $e$, some students from $S^{1}$ or $S^{2}$ are matched with $c_{j}$ and $c_{k}$. For each student $p_{i}$, her preference list contains course $c_{j}$ only:
$p_{j}:c_{j}$
The preference list of every $c_{j}\in C^{\prime}$ depends on the vertices adjacent to $u_{j}$. For every $w_{i}$ adjacent to $u_{j}$, $c_{j}$ includes in its preference list either $s_{i}^{1}$ or $s_{i}^{2}$, as well as $s_{i}^{3}$. Additionally, it includes student $p_{j}$. The order of all students is derived from the master list. If, for example, $u_{j}$ is adjacent to $w_{3}$, $w_{4}$ and $w_{5}$, the preference list of $c_{j}$ is given below. Here the superscript $l_{i}$ indicates a number that can be either $1$ or $2$. If $u_{j}$ is matched in $G$ then $c_{j}$ is matched with some agent from $S^{1}$, $S^{2}$ or $S^{3}$, otherwise it is matched with $p_{j}$.
$c_{j}:s_{3}^{l_{3}}\succ_{c_{j}}s_{4}^{l_{4}}\succ_{c_{j}}s_{5}^{l_{5}}\succ_{c_{j}}s_{3}^{3}\succ_{c_{j}}s_{4}^{3}\succ_{c_{j}}s_{5}^{3}\succ_{c_{j}}p_{j}$
The preference lists of every $d_{i}^{1}\in D^{1}$, $d_{i}^{2}\in D^{2}$, and $d_{i}^{3}\in D^{3}$ are:
$d_{i}^{1}$ $:s_{i}^{1}\succ_{d_{i}^{1}}s_{i}^{2}$
$d_{i}^{2}$ $:s_{i}^{2}\succ_{d_{i}^{2}}s_{i}^{3}$
$d_{i}^{3}$ $:s_{i}^{3}\succ_{d_{i}^{3}}s_{i}^{4}$
Finally, the preference list of $e$ contains every student from $S^{3}$ in increasing subscript order. The students from $S^{3}$ matched to $e$ correspond to vertices in $G$ that are unmatched.
$e:s_{1}^{3}\succ_{e}s_{2}^{3}\succ_{e}...\succ_{e}s_{n_{2}}^{3}$
An example of this transformation is shown in Figure 6.
$u_{1}$$u_{2}$$w_{1}$$w_{2}$$w_{3}$
$s_{1}^{1}$ $:c_{1}\succ_{s_{1}^{1}}d_{1}^{1}$
$s_{1}^{2}$ $:c_{2}\succ_{s_{1}^{2}}d_{1}^{1}\succ_{s_{1}^{2}}d_{1}^{2}$
$s_{1}^{3}$ $:d_{1}^{2}\succ_{s_{1}^{3}}d_{1}^{3}\succ_{s_{1}^{3}}c_{j}\succ_{s_{1}^{3}}c_{k}\succ_{s_{1}^{3}}e$
$s_{1}^{4}$ $:d_{1}^{3}$
$c_{1}$ $:s_{1}^{1}\succ_{c_{1}}s_{2}^{1}\succ_{c_{1}}s_{3}^{1}\succ_{c_{1}}s_{1}^{3}\succ_{c_{1}}s_{2}^{3}\succ_{c_{1}}s_{3}^{3}\succ_{c_{1}}p_{1}$
$e$ $:s_{1}^{3}\succ_{e}s_{2}^{3}\succ_{e}s_{3}^{3}$
Figure 6: Example of a Min-MM instance, with $K=2$, and some of the preference lists in the transformed instance of ca-ml
We first show that a maximal matching $M$ of size at most $K$ in $G$ implies a course-complete first-coalition-stable matching $M^{\prime}$ in $I$. For every pair $(u_{j},w_{i})\in M$, we add to $M^{\prime}$ either the pair $(s_{i}^{1},c_{j})$ or the pair $(s_{i}^{2},c_{j})$, depending on whether $s_{i}^{1}$ or $s_{i}^{2}$ has $c_{j}$ in her preference list. If $(s_{i}^{1},c_{j})\in M^{\prime}$, then we add $(s_{i}^{2},d_{i}^{1})$ to $M^{\prime}$, otherwise we add $(s_{i}^{1},d_{i}^{1})$. We also add $(s_{i}^{3},d_{i}^{2})$ and $(s_{i}^{4},d_{i}^{3})$ to $M^{\prime}$. If $w_{i}$ has no partner in $M$, we add the following pairs to $M^{\prime}$: $(s_{i}^{1},d_{i}^{1})$, $(s_{i}^{2},d_{i}^{2})$, $(s_{i}^{3},e)$, and $(s_{i}^{4},d_{i}^{3})$. Finally, if vertex $u_{j}$ is not matched in $M$, add $(p_{j},c_{j})$ to $M^{\prime}$. The matching $M^{\prime}$ is course-complete, as every course in $C^{\prime}$ or $D$ is matched with one student, and $e$ is matched with $n_{2}-K$ students. It is also first-coalition-stable, as we show below.
Lemma 2.
$M^{\prime}$ is a first-coalition-stable matching.
Proof.
We prove the first-coalition stability of $M^{\prime}$ by showing that no student can belong to a first-blocking coalition. Every $s_{i}^{1}\in S^{1}$ is matched with either some $c_{j}\in C^{\prime}$, which is her most preferred course, or with $d_{i}^{1}$. As $O(d_{i}^{1})>O(c_{j})$, $(s_{i}^{1},\{c_{j}\})$ cannot form a first-blocking coalition. Similarly, every $s_{i}^{2}\in S^{2}$ is matched with either some $c_{k}\in C^{\prime}$, her most preferred course, with $d_{i}^{1}$, or with $d_{i}^{2}$. If $s_{i}^{2}$ is matched with $c_{k}$ or with $d_{i}^{1}$, the same reasoning applies, and if she is matched with $d_{i}^{2}$, then $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$, so there is no blocking coalition containing $s_{i}^{2}$. Every $s_{i}^{4}\in S^{4}$ is matched with the only course in her preference list, $d_{i}^{3}$.
We now consider $s_{i}^{3}\in S^{3}$. This student is matched with either $d_{i}^{2}$ or with $e$. In the first case, she gets her most preferred course. In the second case, $(s_{i}^{2},d_{i}^{2})\in M$ and, because $M$ is maximal in $G$, $u_{j}$ and $u_{k}$ are matched with some $w_{l}$ and $w_{m}$. Thus, there exist $s_{l}^{1}$ (or $s_{l}^{2}$) and $s_{m}^{1}$ (or $s_{m}^{2}$) such that $c_{j}$ and $c_{k}$ are matched with them. Thus, $s_{i}^{3}$ could form a first-blocking coalition with $d_{i}^{3}$ only. However, $O(d_{i}^{3})<O(e)$, so $(s_{i}^{3},\{d_{i}^{3}\})$ is not a first-blocking coalition. Finally, every $p_{j}\in P$ is either matched with $c_{j}$, the only course in her preference list, or she has no partner. In this second case, $c_{j}$ is matched with some $s_{i}^{1}$ or $s_{i}^{2}$, so $c_{j}$ prefers its current student over $p_{j}$, and thus $p_{j}$ cannot be part of a first-blocking coalition. ∎
Now, assume that there exists a course-complete first-coalition-stable matching $M^{\prime}$ in $I$. We now show how to obtain a maximal matching $M$ of size at most $K$ in $G$. First, as $M^{\prime}$ is course-complete, $|S_{M^{\prime}}(e)|=q^{+}(e)=n_{2}-K$, so there are $n_{2}-K$ students $s_{i}^{3}\in S^{3}$ matched to $e$. For each pair $(s_{i}^{3},e)\in M^{\prime}$, $s_{i}^{3}$ would prefer to be matched with $d_{i}^{2}$. Thus, because $M^{\prime}$ is first-coalition-stable, we must have that $(s_{i}^{2},d_{i}^{2})\in M^{\prime}$, which in turn implies that $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$. Likewise, $s_{i}^{3}$ prefers to be matched with $d_{i}^{3}$ and with either $c_{j}$ or with $c_{k}$. Hence, because $M^{\prime}$ is first-coalition-stable, and since $d_{i}^{3}$ ranks $s_{i}^{3}$ first, both $c_{j}$ and $c_{k}$ must have better students than $s_{i}^{3}$, so students from $S^{1}$ or $S^{2}$ (or possibly $S^{3}$). From this we get that, due to $M^{\prime}$ being course-complete, $(s_{i}^{4},d_{i}^{3})\in M^{\prime}$. Thus, if $(s_{i}^{3},e)\in M^{\prime}$, no student from $S_{i}$ is matched with a course from $C^{\prime}$. These $S_{i}$ subsets correspond to $n_{2}-K$ vertices $w_{i}\in G$ that are not matched in the maximal matching $M$.
Let us now consider the $K$ remaining $S_{i}$ subsets. For some of these subsets, there is no $s_{i}^{l}\in S_{i}$ and $c_{j}\in C^{\prime}$ such that $(s_{i}^{l},c_{j})\in M^{\prime}$. In these cases, we have that $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$, $(s_{i}^{2},d_{i}^{2})\in M^{\prime}$ and $(s_{i}^{3},d_{i}^{3})\in M^{\prime}$, as $M^{\prime}$ is first-coalition-stable and course-complete, so every $d_{i}^{l}\in D_{i}$ must have one student. In these cases, we have that both $c_{j}$ and $c_{k}$ are matched with students from $S^{1}$, $S^{2}$ or $S^{3}$, to prevent $s_{i}^{3}$ forming a first-blocking coalition with either $c_{j}$ or $c_{k}$. For the remaining subsets $S_{i}$, there exists exactly one pair of the form $(s_{i}^{l},c_{j})$ in $M^{\prime}$.
Lemma 3.
If $(s_{i}^{l},c_{j})\in M^{\prime}$ for some $s_{i}^{l}\in S_{i}$, then $M^{\prime}$ includes no other pair between a member of $S_{i}$ and a member of $C^{\prime}$.
Proof.
If $s_{i}^{1}$ is matched with $c_{j}$, then $(s_{i}^{2},d_{i}^{1})$ must be in $M^{\prime}$, otherwise $d_{i}^{1}$ has no students and is not course-complete. Thus, $(s_{i}^{3},d_{i}^{2})\in M^{\prime}$ and $(s_{i}^{4},d_{i}^{3})\in M^{\prime}$. If $s_{i}^{2}$ is matched with $c_{k}$, then, by the same reasoning, $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$, $(s_{i}^{3},d_{i}^{2})\in M^{\prime}$ and $(s_{i}^{4},d_{i}^{3})\in M^{\prime}$. Finally, if $(s_{i}^{3},c_{j})\in M^{\prime}$ (or $(s_{i}^{3},c_{k})\in M^{\prime}$), then, to preserve first-coalition stability, $(s_{i}^{3},d_{i}^{3})\in M^{\prime}$, $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$, and $(s_{i}^{2},d_{i}^{2})\in M^{\prime}$. ∎
Hence, for every $S_{i}\subseteq S$, either one $s_{i}^{l}\in S_{i}$ is matched with some $c_{j}$ or none are. If some $s_{i}^{l}\in S_{i}$ is matched with some $c_{j}$, add $(u_{j},w_{i})$ to the matching $M$ in $G$, which is a valid matching because every vertex is matched with at most one, adjacent, vertex. As there are at least $n_{2}-K$ subsets where no member is matched with some $c_{j}$, we have that there are at most $K$ $S_{i}$ subsets where exactly one member is matched with one $c_{j}$. This means that the matching $M$ is of size at most $K$. We now show that $M$ is maximal.
Lemma 4.
The matching $M$ in $G$, of size at most $K$, is maximal.
Proof.
Looking for a contradiction, assume that it is not, so that there exists unmatched $u_{j},w_{i}$ that are adjacent to each other. This implies that in $M^{\prime}$, $s_{i}^{3}$ is matched with either $e$ or $d_{i}^{3}$. In the first case, $c_{j}$ is matched with $p_{j}$, and $d_{i}^{3}$ is matched with $s_{i}^{4}$. However, $s_{i}^{3}$ prefers $\{d_{i}^{3},c_{j}\}$ over $e$, $c_{j}$ prefers $s_{i}^{3}$ over $p_{j}$, and $d_{i}^{3}$ prefers $s_{i}^{3}$ over $s_{i}^{4}$. Thus, $(s_{i}^{3},\{d_{i}^{3},c_{j}\})$ is a first-blocking coalition.
In the second case, $c_{j}$ is matched with $p_{j}$. However, $s_{i}^{3}$ prefers $c_{j}$ over being unmatched, as $O(d_{i}^{3})=1$ and $T(s_{i}^{3})=2$, and $c_{j}$ prefers $s_{i}^{3}$ over $p_{j}$. Thus, $(s_{i}^{3},\{c_{j}\})$ is a first-blocking coalition. In either case we get a contradiction to the stability of $M^{\prime}$. Therefore, $M$ is maximal. ∎
Hence, there exists a maximal matching of size at most $K$ in $G$ if and only if there exists a course-complete first-coalition-stable matching in $I$. As Min-MM is NP-complete, so is determining the existence of a course-complete first-coalition-stable matching in a ca-ml instance. ∎
In fact, it is easy to see that the proof above applies to coalition stability as well. Hence, we obtain the following result.
Corollary 1.
Determining whether a ca-ml instance has a course-complete coalition-stable matching is NP-hard, even if every student takes up to $2$ credits, every course has $1$ or $2$ credits, and each student’s preference list is of length at most $5$.
The main difference between this corollary and Theorem 5 is that in Theorem 5, courses are also ordered in a master list and every course has upper quota $1$, but the length of students’ preference lists is not bounded.
As an aside, one can determine whether a ca-ml instance $I$ has a student-complete first-coalition-stable matching in polynomial time, as we now show.
Proposition 3.
Determining whether a ca-ml instance $I$ has a student-complete first-coalition-stable (feasible) matching is solvable in polynomial time.
Proof.
In order to find a student-complete first-coalition-stable matching, or show that none exists, we use a variant of Serial Dictatorship. In this variant, starting from the beginning of the preference list, for each student $s_{i}$ we attempt to find a subset $C_{i}\subseteq P(s_{i})$ such that $s_{i}$ can be matched with $C_{i}$ (i.e. the resulting matching is first-coalition stable and no constraints are violated) and $O(C_{i})=T(s_{i})$. If such a $C_{i}$ exists, then it is unique. Assume the opposite, so that $s_{i}$ can be matched with courses from subsets $C_{i}\subseteq P(s_{i})$ and $C_{i}^{\prime}\subseteq P(s_{i})$, such that $C_{i}\neq D_{i}$, $O(C_{i})=O(D_{i})=T(s_{i})$, both $C_{i}$ and $D_{i}$ are feasible, and every course in both $C_{i}$ and $D_{i}$ has space for $s_{i}$. Let $c_{i}$ be $s_{i}$’s most preferred course in $C_{i}\backslash D_{i}$ and $d_{i}$ her most preferred course in $D_{i}\backslash C_{i}$. Without loss of generality, assume that $c_{i}\succ_{s_{i}}d_{i}$. Then $s_{i}$ prefers $C_{i}$ over $D_{i}$, and so if matched with the courses in $D_{i}$ she will form a first-blocking coalition with a subset of courses of $C_{i}$. Hence there is at most one possible $C_{i}$ that satisfies all conditions.
The algorithm proceeds as follows. First, the first student in the master list, $s_{1}$, selects the unique subset $C_{1}$, should it exist, and they are matched together. Then $s_{2}$ selects the unique subset $C_{2}$, and so on. If for some $s_{i}$ such a $C_{i}$ does not exist, then no student-complete first-coalition-stable matching exists, because $s_{i}$ cannot choose a subset of courses that allows her to take her maximum number of credits, and the previous students in the master list cannot choose different courses, which may have freed some space in some $c_{j}\in P(s_{i})$ and allowed $s_{i}$ to select a different subset. Otherwise, for every $s_{i}$, $O(C_{i})=T(s_{i})$, so we have found a student-complete first-coalition-stable matching. ∎
We now show that determining the existence of a student-complete (or course-complete) pair-size-stable matching is NP-hard. In order to prove this result, we will reduce from Exact-MM. In this problem, we have a graph $G$ and a positive integer $K$, and the goal is to determine whether there exists a maximal matching of size exactly $K$. This problem is NP-complete, even if $G$ is a subdivision graph iwama_2002_stable ; manlove_2002_hard (manlove_algorithmics_2013, , page 133), which implies that it is bipartite and all vertices in one of the sides have degree $2$.
Theorem 8.
It is NP-hard to determine whether an instance of ca-ml has a student-complete (or course-complete) pair-size-stable matching, even if both students and courses are ordered according to master lists, every student takes up to $2$ credits, and every course has $1$ or $2$ credits.
Proof.
We reduce from Exact-MM, which is NP-complete, even if $G$ is bipartite and all vertices in one of the sides have degree $2$. Let $G=(U,W,E)$ be a graph representing an instance of Exact-MM, where $U$ and $W$ are the two sets of vertices, $|U|=n_{1}$ and $|W|=n_{2}$. Every vertex in $W$ has degree $2$. Our goal is to find a maximal matching of size $K$. We build an instance $I$ of ca-ml, with students $S$ and courses $C$, as follows.
For every $w_{i}\in W$, we create the following students: $s_{i}^{1}$, $s_{i}^{2}$, and $s_{i}^{3}$, with $T(s_{i}^{1})=T(s_{i}^{2})=T(s_{i}^{3})=2$. We also create courses $d_{i}^{1}$ and $d_{i}^{2}$, with $O(d_{i}^{1})=O(d_{i}^{2})=2$. Students $s_{i}^{1}$, $s_{i}^{2}$ and $s_{i}^{3}$ belong to subset $S_{i}\subseteq S$, and courses $d_{i}^{1}$ and $d_{i}^{2}$ belong to subset $D_{i}\subseteq C$. For every $w_{i}\in W$, $s_{i}^{1}\in S^{1}\subseteq S$, $s_{i}^{2}\in S^{2}\subseteq S$, and $s_{i}^{3}\in S^{3}\subseteq S$. Similarly, $d_{i}^{1}\in D^{1}\subseteq C$, and $d_{i}^{2}\in D^{2}\subseteq C$.
For every $u_{j}\in U$, we create a course $c_{j}$, with $O(c_{j})=1$. Additionally, we create three courses $e_{1}$, $e_{2}$ and $f$, with $O(e_{1})=O(e_{2})=2$ and $O(f)=1$, as well as $n_{1}-K$ students labelled $p_{1}$ to $p_{n_{1}-K}$, such that, for every $p_{k}$, $T(p_{k})=1$. The upper quota of every course is $1$, except for $e_{1}$ and $e_{2}$, which have upper quota $n_{2}-K$, and $f$, which has upper quota $K$. For every $u_{j}\in U$, $c_{j}\in C^{\prime}\subseteq C$. Every $p_{k}$ belongs to $P\subseteq S$.
The master list of students (MLS) is
$s_{1}^{1}\succ_{MLS}s_{1}^{2}\succ_{MLS}s_{2}^{1}\succ_{MLS}s_{2}^{2}\succ_{MLS}...\succ_{MLS}s_{n_{2}}^{1}\succ_{MLS}s_{n_{2}}^{2}\succ_{MLS}\\ \succ_{MLS}s_{1}^{3}\succ_{MLS}s_{2}^{3}\succ_{MLS}...\succ_{MLS}s_{n_{2}}^{3}\succ_{MLS}p_{1}\succ_{MLS}p_{2}\succ_{MLS}...\succ_{MLS}p_{n_{1}-K}$ (2)
That is, the first students belong to subsets $S^{1}$ and $S^{2}$, such that every $s_{i}^{1}$ is followed by $s_{i}^{2}$. Then we place students from $S^{3}$, and finally students from $P$, in all cases in increasing subscript order. The master list of courses (MLC) is
$c_{1}\succ_{MLC}c_{2}\succ_{MLC}...\succ_{MLC}c_{n_{1}}\succ_{MLC}d_{1}^{1}\succ_{MLC}d_{1}^{2}\succ_{MLC}d_{2}^{1}\succ_{MLC}d_{2}^{2}\succ_{MLC}...\succ_{MLC}\\ \succ_{MLC}d_{n_{2}}^{1}\succ_{MLC}d_{n^{2}}^{2}\succ_{MLC}e_{1}\succ_{MLC}e_{2}\succ_{MLC}f$ (3)
That is, the first courses belong to $C^{\prime}$, then the next courses are from $D^{1}$ and $D^{2}$, such that for each $w_{i}\in W$, $d_{i}^{1}\succ_{MLC}d_{i}^{2}$, and finally we have courses $e_{1}$, $e_{2}$ and $f$.
Let $w_{i}\in W$ be a vertex from $G$, adjacent to $u_{j}$ and $u_{k}$ (with $j<k$). We build the preference lists of $s_{i}^{1}$, $s_{i}^{2}$ and $s_{i}^{3}$ as follows.
$s_{i}^{1}$ $:c_{j}\succ_{s_{i}^{1}}d_{i}^{1}\succ_{s_{i}^{1}}f$
$s_{i}^{2}$ $:c_{k}\succ_{s_{i}^{2}}d_{i}^{1}\succ_{s_{i}^{2}}d_{i}^{2}\succ_{s_{i}^{2}}f$
$s_{i}^{3}$ $:c_{j}\succ_{s_{i}^{3}}c_{k}\succ_{s_{i}^{3}}d_{i}^{2}\succ_{s_{i}^{3}}e_{1}\succ_{s_{i}^{3}}e_{2}$
If in the Exact-MM instance $w_{i}$ is matched with $u_{j}$, then $s_{i}^{1}$ is to be matched with $c_{j}$, $s_{i}^{2}$ with $d_{i}^{1}$ and $s_{i}^{3}$ with $d_{i}^{2}$. If, however, $w_{i}$ is unmatched, $s_{i}^{1}$ is to be matched with $d_{i}^{1}$, $s_{i}^{2}$ with $d_{i}^{2}$, and $s_{i}^{3}$ with $e_{1}$ and $e_{2}$. Student $s_{i}^{3}$ contains courses $c_{j}$ and $c_{k}$ in her preference list to make sure that a course-complete first-coalition-stable matching in $I$ corresponds to a maximal matching in $G$, so that if $s_{i}^{3}$ is matched with $e_{1}$ and $e_{2}$, some students from $S^{1}$ or $S^{2}$ are matched with $c_{j}$ and $c_{k}$. For each student $p_{k}$, her preference list contains every course in $C^{\prime}$ in increasing subscript order:
$p_{k}:c_{1}\succ_{p_{k}}c_{2}\succ_{p_{k}}...\succ_{p_{k}}c_{n_{1}}$
The preference list of every $c_{j}\in C^{\prime}$ depends on the vertices adjacent to $u_{j}$. For every $w_{i}$ adjacent to $u_{j}$, $c_{j}$ includes in its preference list either $s_{i}^{1}$ or $s_{i}^{2}$, as well as $s_{i}^{3}$. Additionally, it includes every student in $P$. The order of all students is derived from the master list. If, for example, $u_{j}$ is adjacent to $w_{3}$, $w_{4}$ and $w_{5}$, the preference list of $c_{j}$ is given below. Here the superscript $l_{i}$ indicates a number that can be either $1$ or $2$. If $u_{j}$ is matched in $G$ then $c_{j}$ is matched with some agent from $S^{1}$ or $S^{2}$, otherwise it is matched with some $p_{k}$.
$c_{j}:s_{3}^{l_{3}}\succ_{c_{j}}s_{4}^{l_{4}}\succ_{c_{j}}s_{5}^{l_{5}}\succ_{c_{j}}s_{3}^{3}\succ_{c_{j}}s_{4}^{3}\succ_{c_{j}}s_{5}^{3}\succ_{c_{j}}p_{1}\succ_{c_{j}}p_{2}\succ_{c_{j}}...\succ_{c_{j}}p_{n_{1}}$
The preference lists of every $d_{i}^{1}\in D^{1}$ and $d_{i}^{2}\in D^{2}$ are:
$d_{i}^{1}$ $:s_{i}^{1}\succ_{d_{i}^{1}}s_{i}^{2}$
$d_{i}^{2}$ $:s_{i}^{2}\succ_{d_{i}^{2}}s_{i}^{3}$
Finally, the preference lists of $e_{1}$ and $e_{2}$ contain every student from $S^{3}$ and the preference list of $f$ contains every student from $S^{1}$ and $S^{2}$, in both cases in increasing subscript order. The students matched to $e_{1}$ and $e_{2}$ correspond to vertices in $G$ that are unmatched, while the students matched to $f$ correspond to vertices that are matched.
$e_{1}$ $:s_{1}^{3}\succ_{e_{1}}s_{2}^{3}\succ_{e_{1}}...\succ_{e_{1}}s_{n_{2}}^{3}$
$e_{2}$ $:s_{1}^{3}\succ_{e_{2}}s_{2}^{3}\succ_{e_{2}}...\succ_{e_{2}}s_{n_{2}}^{3}$
$f$ $:s_{1}^{1}\succ_{f}s_{1}^{2}\succ_{f}s_{2}^{1}\succ_{f}s_{2}^{2}\succ_{f}...\succ_{f}s_{n_{2}}^{1}\succ_{f}s_{n_{2}}^{2}$
An example of this transformation is shown in Figure 7.
$u_{1}$$u_{2}$$w_{1}$$w_{2}$$w_{3}$
$s_{1}^{1}$ $:c_{1}\succ_{s_{1}^{1}}d_{1}^{1}\succ_{s_{1}^{1}}f$
$s_{1}^{2}$ $:c_{2}\succ_{s_{1}^{2}}d_{1}^{1}\succ_{s_{1}^{2}}d_{1}^{2}\succ_{s_{1}^{2}}f$
$s_{1}^{3}$ $:c_{1}\succ_{s_{1}^{3}}c_{2}\succ_{s_{1}^{3}}d_{1}^{2}\succ_{s_{1}^{3}}e_{1}\succ_{s_{1}^{3}}e_{2}$
$c_{1}$ $:s_{1}^{1}\succ_{c_{1}}s_{2}^{1}\succ_{c_{1}}s_{3}^{1}\succ_{c_{1}}s_{1}^{3}\succ_{c_{1}}s_{2}^{3}\succ_{c_{1}}s_{3}^{3}\succ_{c_{1}}p_{1}$
$e_{1}$ $:s_{1}^{3}\succ_{e_{1}}s_{2}^{3}\succ_{e_{1}}s_{3}^{3}$
$f$ $:s_{1}^{1}\succ_{f}s_{1}^{2}\succ_{f}s_{2}^{1}\succ_{f}s_{2}^{2}\succ_{f}s_{3}^{1}\succ_{f}s_{3}^{2}$
Figure 7: Example of an Exact-MM instance, with $K=2$, and some of the preference lists in the transformed instance of ca-ml
We first show that a course-complete matching in $I$ is student-complete, and vice versa.
Lemma 5.
Let $M^{\prime}$ be a matching in $I$. Then $M^{\prime}$ is course-complete if and only if it is student-complete.
Proof.
There are $n_{1}$ courses in $C^{\prime}$, with upper quota $1$ and number of credits $1$. There are also $n_{2}$ courses in each one of $D^{1}$ and in $D^{2}$, with upper quota $1$ and number of credits $2$. Finally, we have courses $e_{1}$ and $e_{2}$, with upper quota $n_{2}-K$ and number of credits $1$, and $f$, with upper quota $K$ and number of credits $1$. Hence, a course-complete matching has $n_{1}\cdot 1\cdot 1+2\cdot n_{2}\cdot 2\cdot 1+2\cdot(n_{2}-K)\cdot 1+1\cdot K\cdot 1=n_{1}+6n_{2}-K$ credits.
Likewise, there are $n_{2}$ students in each one of $S^{1}$, $S^{2}$ and $S^{3}$, with number of credits $2$, and $n_{1}-K$ students in $P$, with number of credits $1$. Thus, a student-complete matching has $3\cdot n_{2}\cdot 2+(n_{1}-K)\cdot 1=n_{1}+6n_{2}-K$ number of credits.
Thus, a course-complete matching has the same number of credits as a student-complete matching. ∎
We now show that a maximal matching $M$ of size $K$ in $G$ implies a student-complete (course-complete) pair-size-stable matching $M^{\prime}$ in $I$. For every pair $(u_{j},w_{i})\in M$, we add to $M^{\prime}$ either the pairs $(s_{i}^{1},c_{j})$ and $(s_{i}^{1},f)$, or the pairs $(s_{i}^{2},c_{j})$ and $(s_{i}^{2},f)$, depending on whether $s_{i}^{1}$ or $s_{i}^{2}$ has $c_{j}$ in her preference list. If $(s_{i}^{1},c_{j})\in M^{\prime}$, then we add $(s_{i}^{2},d_{i}^{1})$ to $M^{\prime}$, otherwise we add $(s_{i}^{1},d_{i}^{1})$. We also add $(s_{i}^{3},d_{i}^{2})$ to $M^{\prime}$. If $w_{i}$ has no partner in $M$, we add the following pairs to $M^{\prime}$: $(s_{i}^{1},d_{i}^{1})$, $(s_{i}^{2},d_{i}^{2})$, $(s_{i}^{3},e_{1})$, and $(s_{i}^{3},e_{2})$. Finally, let $j_{k}$ ($1\leq k\leq n_{1}-K$) be a sequence such that $w_{j_{k}}$ is unmatched in $M$. Then add $(p_{k},w_{j_{k}})$ to $M^{\prime}$. The matching $M^{\prime}$ is course-complete, as every course in $C^{\prime}$, $D^{1}$ and $D^{2}$ is matched with one student, $e_{1}$ and $e_{2}$ are matched with $n_{2}-K$ students, and $f$ is matched with $K$ students. It is also student-complete, because for every student $s\in S$, $T(s)=O(C_{M^{\prime}}(s))$. We show below that it is pair-size-stable.
Lemma 6.
$M^{\prime}$ is a pair-size-stable matching.
Proof.
We prove the pair-size stability of $M^{\prime}$ by showing that no student can belong to a size-blocking pair. Every $s_{i}^{1}$ is matched with either $f$ and some $c_{j}$, which is her most preferred course, or with $d_{i}^{1}$. In either case, there is no course with which $s_{i}^{1}$ can form a size-blocking pair. Similarly, every $s_{i}^{2}$ is matched with either $f$ and some $c_{k}$, her most preferred course, with $d_{i}^{1}$, or with $d_{i}^{2}$. If $s_{i}^{2}$ is matched with $f$ and $c_{k}$, or with $d_{i}^{1}$, the same reasoning applies, and if she is matched with $d_{i}^{2}$, then $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$, so there is no size-blocking pair containing $s_{i}^{2}$.
We now consider $s_{i}^{3}$. This student is matched with either $d_{i}^{2}$ or with $e_{1}$ and $e_{2}$. In the first case, there is no course with an equal number of credits that $s_{i}^{3}$ prefers more, so there is no possible size-blocking pair. In the second case, $(s_{i}^{2},d_{i}^{2})\in M^{\prime}$ and, because $M$ is maximal in $G$, $u_{j}$ and $u_{k}$ are matched with some $w_{l}$ and $w_{m}$. Thus, there exist $s_{l}^{1}$ (or $s_{l}^{2}$) and $s_{m}^{1}$ (or $s_{m}^{2}$) such that $c_{j}$ and $c_{k}$ are matched with them. Hence, $s_{i}^{3}$ cannot form a size-blocking pair with any other course. Finally, every $p_{k}\in P$ is matched with some $c_{j}\in C^{\prime}$. Let $c_{l}$ be a course that $p_{k}$ prefers over $c_{j}$. Then $c_{l}$ is matched with either some $s_{i}^{1}$ or $s_{i}^{2}$, or with a $p_{k}^{\prime}$ such that $p_{k}^{\prime}\succ_{MLS}p_{k}$. Thus, $p_{j}$ cannot form a size-blocking pair with $c_{l}$. ∎
Now, assume that there exists a course-complete (student-complete) pair-size-stable matching $M^{\prime}$ in $I$. We now show how to get a maximal matching $M$ of size $K$ in $G$. First, as $M^{\prime}$ is student-complete, every $p_{k}\in P$ is matched with some $c_{j}\in C^{\prime}$. As $M^{\prime}$ is course-complete, the $K$ remaining courses in $C^{\prime}$ are matched with students from $S^{1}$, $S^{2}$, or $S^{3}$. Similarly, as $M^{\prime}$ is course-complete, $|S_{M}(f)|=q^{+}(f)$, that is, course $f$ fills its capacity. Every student who has $f$ in her preference list belongs to either $S^{1}$ or $S^{2}$, and because $M^{\prime}$ is student-complete every student matched to $f$ must be matched to some $c_{j}$, so the $K$ courses from $C^{\prime}$ that are not matched with students from $P$ are matched with students from $S^{1}$ and $S^{2}$. We now show that for each subset $S_{i}\subseteq S$, either $s_{i}^{1}$ or $s_{i}^{2}$ is matched with some $c_{j}\in C^{\prime}$.
Lemma 7.
If $(s_{i}^{1},c_{j})\in M^{\prime}$ or $(s_{i}^{2},c_{j})\in M^{\prime}$, then $M^{\prime}$ includes no other pair between a member of $S_{i}$ and a member of $C^{\prime}$.
Proof.
If $s_{i}^{1}$ is matched with $c_{j}$, then $(s_{i}^{2},d_{i}^{1})$ must be in $M^{\prime}$, otherwise $d_{i}^{1}$ has no students and is not course-complete. Thus, $(s_{i}^{3},d_{i}^{2})\in M^{\prime}$, or else $d_{i}^{2}$ has no students. If $s_{i}^{2}$ is matched with $c_{k}$, then, by the same reasoning, $(s_{i}^{2},d_{i}^{1})\in M^{\prime}$ and $(s_{i}^{3},d_{i}^{2})\in M^{\prime}$.∎
Thus, we can build a matching $M$ in $G$ of size $K$ by taking every pair of the form $(s_{i}^{l},c_{j})\in M$, of which there are $K$, and adding the pair $(u_{j},w_{i})$ to $M$. For every $S_{i}$ such that no $s_{i}^{l}\in S_{i}$ is matched to some $c_{j}\in C^{\prime}$, we have that $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$, $(s_{i}^{2},d_{i}^{2})\in M^{\prime}$, $(s_{i}^{3},e_{1})\in M^{\prime}$, and $(s_{i}^{3},e_{2})\in M^{\prime}$. We now show that $M$ is a maximal matching in $G$.
Lemma 8.
The matching $M$ in $G$, of size $K$, is maximal.
Proof.
Looking for a contradiction, assume that it is not, so that there exists unmatched $u_{j},w_{i}$ that are adjacent to each other. This implies that in $M^{\prime}$, $c_{j}$ is matched with some $p_{k}$ and $s_{i}^{3}$ is matched with $e_{1}$ and $e_{2}$. However, $s_{i}^{3}$ prefers $c_{j}$ over $e_{1}$ and $e_{2}$ and $c_{j}$ prefers $s_{i}^{3}$ over $p_{k}$. Thus, $(s_{i}^{3},c_{j})$ is a size-blocking pair, a contradiction to the stability of $M^{\prime}$. Therefore, $M$ is maximal. ∎
Hence, there exists a maximal matching of size $K$ in $G$ if and only if there exists a course-complete (student-complete) pair-size-stable matching in $I$. As Exact-MM is NP-complete, determining the existence of a course-complete (student-complete) pair-size-stable matching is NP-hard. ∎
We can modify this transformation to show that determining the existence of a course-complete (but not necessarily student-complete) pair-size-stable matching is NP-hard even if all conditions from the theorem statement hold and the length of the preference list of every student is bounded. We do so by removing every agent in $P$, and creating a new $p_{j}$ agent for every $c_{j}\in C^{\prime}$. The preference list of this $p_{j}$ contains $c_{j}$ only, and these $p_{j}$’s are at the end of the master list of students.
5.2 With lower quotas
Until now we have assumed that courses do not have a minimum number of students that they require in order to run, which are represented by lower quotas. However, this is a very natural requirement, present in many universities, and so we now define an extension to the Course Allocation problem involving lower quotas. A course $c_{j}$ is said to have a lower quota, denoted by $q^{-}(c_{j})$, if in a matching $M$ the number of students allocated to $c_{j}$ must be at least the lower quota, i.e., $|S_{M}(c_{j})|\geq q^{-}(c_{j})$. When courses have lower quotas, there are two variations to consider. In the first one, every course in the Course Allocation instance must have a number of students at least as large as its lower quota. The goal is to find a stable matching (where every course satisfies its lower and upper quota) for any given definition of stability. This problem is called ca-x-lq-nc-find, where x denotes the definition of stability. In the case of ca-ml, the problem is called ca-x-ml-lq-nc-find.
The second variation models the scenario where in a given matching $M$ if course $c_{j}$ does not have at least as many students as its lower quota then it may be closed, and no students are matched to it. In this case an empty matching, where every course is closed, is trivially stable, so our goal is to find a maximum size stable matching. This problem is called ca-x-lq-cl-max, where x denotes the definition of stability. In the case of ca-ml, the problem is called ca-x-ml-lq-cl-max. In this variation we do not allow for the existence of a coalition of several students and a closed course that would demand for this course to stay open (as in biro_college_2010 ), because in real-life settings lower quotas are usually too high for students to coordinate in such a fashion.
From Theorems 7 and 8 and Corollary 1 we immediately get these two results.
Observation 4.
Given an instance $I$ of ca-ml with lower quotas, ca-ps-ml-lq-nc-find, ca-c-ml-lq-nc-find and ca-fc-ml-lq-nc-find are NP-hard, even if every student takes up to $2$ credits, every course has $1$ or $2$ credits, the length of the students’ preference lists is bounded, and there are no downward-feasible constraints.
Proof.
Modify the instances from Theorems 7 and 8 and Corollary 1 so that the lower quota of every course is the same as the upper quota. Then a stable matching respecting lower quotas exists if and only if a course-complete stable matching exists, for all three definitions of stability. ∎
Observation 5.
ca-ps-ml-lq-cl-max, ca-c-ml-lq-cl-max and ca-fc-ml-lq-cl-max are NP-hard, even if every student takes up to $2$ credits, every course has $1$ or $2$ credits, the length of the students’ preference lists is bounded, and there are no downward-feasible constraints.
Proof.
Under pair-size stability, coalition stability or first-coalition stability, finding a maximum size stable matching is hard without lower quotas, so it remains hard with lower quotas. ∎
Similarly, given an instance of ca with lower quotas, determining the existence of a pair-stable matching is NP-complete, which is an immediate consequence of Theorem 3. However, in the case of master lists, we know that every instance of ca-ml has a unique pair-stable matching. Thus, in order to determine whether an instance of ca-ml with lower quotas has a pair-stable matching that respects lower quotas, we just need to run Serial Dictatorship and check whether in the resulting matching every course has enough students to fill its lower quota. Thus, we obtain the following Observation.
Observation 6.
ca-p-ml-lq-nc-find is solvable in polynomial time.
On the other hand, if we allow for course closures, ca-p-ml-lq-cl-max is NP-hard.
Theorem 9.
It is NP-complete to determine if an instance of ca-ml with lower quotas and closures has a student-complete pair-stable matching, even if both students and courses are ordered according to master lists, every student takes up to $2$ credits, every course is of size $1$ or $2$ and with lower quota $1$ or $2$, and there are no downward-feasible constraints.
Proof.
Verifying whether a matching is student-complete and pair-stable, and respects lower quotas, can be done in polynomial time, so the problem belongs to NP. We prove NP-completeness by reducing from Exact-MM, which is NP-complete, even if $G$ is bipartite and all vertices in one of the sides have degree $2$. Let $G=(U,W,E)$ be a graph representing an instance of Exact-MM, where $U$ and $W$ are the two sets of vertices, $|U|=n_{1}$ and $|W|=n_{2}$. Every vertex in $W$ has degree $2$. Our goal is to find a maximal matching of size $K$. We build an instance $I$ of ca-ml-lq-cl, with students $S$ and courses $C$, as follows.
For every $w_{i}\in W$, we create the following students: $s_{i}^{1}$, $s_{i}^{2}$, $s_{i}^{3}$, and $s_{i}^{4}$, with $T(s_{i}^{1})=T(s_{i}^{2})=T(s_{i}^{3})=T(s_{i}^{4})=2$. We also create courses $d_{i}^{1}$ and $d_{i}^{2}$, with $O(d_{i}^{1})=O(d_{i}^{2})=2$, as well as courses $b_{i}^{1}$ and $b_{i}^{2}$, with $O(b_{i}^{1})=O(b_{i}^{2})=1$. Students $s_{i}^{1}$, $s_{i}^{2}$, $s_{i}^{3}$ and $s_{i}^{4}$ belong to subset $S_{i}\subseteq S$, and courses $d_{i}^{1}$, $d_{i}^{2}$ and $d_{i}^{3}$ belong to subset $D_{i}\subseteq C$. For every $w_{i}\in W$, $s_{i}^{1}\in S^{1}\subseteq S$, $s_{i}^{2}\in S^{2}\subseteq S$, $s_{i}^{3}\in S^{3}\subseteq S$, and $s_{i}^{4}\in S^{4}\subseteq S$. Similarly, $d_{i}^{1}\in D^{1}\subseteq C$, $d_{i}^{2}\in D^{2}\subseteq C$ and $b_{i}^{1},b_{i}^{2}\in B\subseteq C$.
For every $u_{j}\in U$, we create a course $c_{j}$, with $O(c_{j})=1$. Additionally, we create two courses $e$ and $f$, with $O(e)=O(f)=2$. The upper quota of every course is $1$, except for $e$, which has upper quota $n_{2}-K$, and every $b_{i}^{l}\in B$, $d_{i}^{1}\in D^{1}$ and $f$, which have upper quota $2$. The lower quota of every course is $1$, except for every $b_{i}^{l}\in B$, every $d_{i}^{1}\in D^{1}$ and $f$, which have lower quota $2$. For every $u_{j}\in U$, $c_{j}\in C^{\prime}\subseteq C$.
We also create $K$ students labelled $p_{1}$ to $p_{K}$ such that, for every $p_{k}$, $T(p_{k})=1$, and two students $q$ and $r$ such that $T(q)=T(r)=2$. For every $p_{k}$, $1\leq k\leq K$, $p_{k}\in P\subseteq S$.
The master list of students (MLS) is
$s_{1}^{1}\succ_{MLS}s_{1}^{4}\succ_{MLS}s_{1}^{2}\succ_{MLS}s_{2}^{1}\succ_{MLS}s_{2}^{4}\succ_{MLS}s_{2}^{2}\succ_{MLS}...\succ_{MLS}s_{n_{2}}^{1}\succ_{MLS}\\ \succ_{MLS}s_{n_{2}}^{4}\succ_{MLS}s_{n_{2}}^{2}\succ_{MLS}s_{1}^{3}\succ_{MLS}s_{2}^{3}\succ_{MLS}...\succ_{MLS}s_{n_{2}}^{3}\succ_{MLS}\\ \succ_{MLS}p_{1}\succ_{MLS}...\succ_{MLS}p_{K}\succ_{MLS}q\succ_{MLS}r$ (4)
That is, the first students belong to subsets $S^{1}$, $S^{4}$ and $S^{2}$, such that every $s_{i}^{1}$ is followed by $s_{i}^{4}$ and then by $s_{i}^{2}$. Then we place students from $S^{3}$, and after them students from $P$, in all cases in increasing subscript order. Finally, we place students $q$ and $r$. The master list of courses (MLC) is
$b_{1}^{1}\succ_{MLC}b_{1}^{2}\succ_{MLC}b_{2}^{1}\succ_{MLC}b_{2}^{2}\succ_{MLC}...\succ_{MLC}b_{n_{2}}^{1}\succ_{MLC}b_{n_{2}}^{2}\succ_{MLC}\\ \succ_{MLC}d_{1}^{1}\succ_{MLC}d_{1}^{2}\succ_{MLC}d_{2}^{1}\succ_{MLC}d_{2}^{2}\succ_{MLC}...\succ_{MLC}d_{n_{2}}^{1}\succ_{MLC}d_{n_{2}}^{2}\succ_{MLC}\\ \succ_{MLC}c_{1}\succ_{MLC}c_{2}\succ_{MLC}...\succ_{MLC}c_{n_{1}}\succ_{MLC}e\succ_{MLC}f$ (5)
That is, the first courses are those from $B$, then the next courses are from $D^{1}$ and $D^{2}$, such that for each $w_{i}\in W$, $d_{i}^{1}\succ_{MLC}d_{i}^{2}$, after them those which belong to $C^{\prime}$, and finally we have courses $e$ and $F$, in all cases in increasing subscript order.
Let $w_{i}\in W$ be a vertex from $G$, adjacent to $u_{j}$ and $u_{k}$ (with $j<k$). We build the preference lists of $s_{i}^{1}$, $s_{i}^{2}$ and $s_{i}^{3}$ as follows.
$s_{i}^{1}$ $:b_{i}^{1}\succ_{s_{i}^{1}}d_{i}^{1}\succ_{s_{i}^{1}}c_{j}$
$s_{i}^{2}$ $:b_{i}^{2}\succ_{s_{i}^{2}}d_{i}^{1}\succ_{s_{i}^{2}}d_{i}^{2}\succ_{s_{i}^{2}}c_{k}$
$s_{i}^{3}$ $:d_{i}^{2}\succ_{s_{i}^{3}}c_{j}\succ_{s_{i}^{3}}c_{k}\succ_{s_{i}^{3}}e$
$s_{i}^{4}$ $:d_{i}^{1}$
If in the Exact-MM instance $w_{i}$ is matched with $u_{j}$, then $s_{i}^{1}$ is to be matched with $c_{j}$ and $b_{i}^{1}$, $s_{i}^{2}$ with $d_{i}^{1}$, $s_{i}^{3}$ with $d_{i}^{2}$, and $s_{i}^{4}$ with $d_{i}^{1}$. If, however, $w_{i}$ is unmatched, $s_{i}^{1}$ is to be matched with $d_{i}^{1}$, $s_{i}^{2}$ with $d_{i}^{2}$, $s_{i}^{3}$ with $e$, and $s_{i}^{4}$ with $d_{i}^{1}$. Student $s_{i}^{3}$ contains courses $c_{j}$ and $c_{k}$ in her preference list to make sure that a course-complete pair-stable matching in $I$ corresponds to a maximal matching in $G$, so that if $s_{i}^{3}$ is matched with $e$, some students from $S^{1}$ or $S^{2}$ are matched with $c_{j}$ and $c_{k}$. The preference lists of agents in $P$ includes every course in $B$.
$p_{k}:b_{1}^{1}\succ_{p_{k}}b_{1}^{2}\succ_{p_{k}}b_{2}^{1}\succ_{p_{k}}b_{2}^{2}\succ_{p_{k}}...\succ_{p_{k}}b_{n_{2}}^{1}\succ_{p_{k}}b_{n_{2}}^{1}$
Agents from $P$ are matched to the courses in $B$ that have some $s_{i}^{l}$ matched to them, which correspond to vertices $w_{i}\in G$ that are matched in $G$. The preference lists of $q$ and $r$ are:
$q$ $:e\succ_{q}f$
$r$ $:f$
The preference list of every $c_{j}\in C^{\prime}$ depends on the vertices adjacent to $u_{j}$. For every $w_{i}$ adjacent to $u_{j}$, $c_{j}$ includes in its preference list either $s_{i}^{1}$ or $s_{i}^{2}$, as well as $s_{i}^{3}$. The order of all students is derived from the master list. If, for example, $u_{j}$ is adjacent to $w_{3}$, $w_{4}$ and $w_{5}$, the preference list of $c_{j}$ is given below. Here the superscript $l_{i}$ indicates a number that can be either $1$ or $2$. If $u_{j}$ is matched in $G$ then $c_{j}$ is matched with some agent from $S^{1}$ or $S^{2}$, otherwise it is unmatched.
$c_{j}:s_{3}^{l_{3}}\succ_{c_{j}}s_{4}^{l_{4}}\succ_{c_{j}}s_{5}^{l_{5}}\succ_{c_{j}}s_{3}^{3}\succ_{c_{j}}s_{4}^{3}\succ_{c_{j}}s_{5}^{3}$
The preference lists of every $b_{i}^{1}$ and $b_{i}^{2}$ in $B$ include $s_{i}^{1}$ in the case of $b_{i}^{1}$, and $s_{i}^{2}$ in the case of $b_{i}^{2}$. Then, we have every student in $P$.
$b_{i}^{1}$ $:s_{i}^{1}\succ_{b_{i}^{1}}p_{1}\succ_{b_{i}^{1}}p_{2}\succ_{b_{i}^{1}}...\succ_{b_{i}^{1}}p_{K}$
$b_{i}^{2}$ $:s_{i}^{2}\succ_{b_{i}^{2}}p_{1}\succ_{b_{i}^{2}}p_{2}\succ_{b_{i}^{2}}...\succ_{b_{i}^{2}}p_{K}$
The preference lists of every $d_{i}^{1}\in D^{1}$, $d_{i}^{2}\in D^{2}$ and of $f$ are:
$d_{i}^{1}$ $:s_{i}^{1}\succ_{d_{i}^{1}}s_{i}^{4}\succ_{d_{i}^{1}}s_{i}^{2}$
$d_{i}^{2}$ $:s_{i}^{2}\succ_{d_{i}^{2}}s_{i}^{3}$
$f$ $:q\succ_{f}r$
Finally, the preference list of $e$ contains every student from $S^{3}$ in increasing subscript order, and then $q$. The students from $S^{3}$ matched to $e$ correspond to vertices in $G$ that are unmatched.
$e:s_{1}^{3}\succ_{e}s_{2}^{3}\succ_{e}...\succ_{e}s_{n_{2}}^{3}\succ_{e}q$
An example of this transformation is shown in Figure 8.
$u_{1}$$u_{2}$$w_{1}$$w_{2}$$w_{3}$
$s_{1}^{1}$ $:b_{1}^{1}\succ_{s_{1}^{1}}d_{1}^{1}\succ_{s_{1}^{1}}c_{1}$
$s_{1}^{2}$ $:b_{1}^{2}\succ_{s_{1}^{2}}d_{1}^{1}\succ_{s_{1}^{2}}d_{1}^{2}\succ_{s_{1}^{2}}c_{2}$
$s_{1}^{3}$ $:d_{1}^{2}\succ_{s_{1}^{3}}c_{1}\succ_{s_{1}^{3}}c_{2}\succ_{s_{1}^{3}}e$
$s_{1}^{4}$ $:d_{1}^{1}$
$p_{1}$ $:b_{1}^{1}\succ_{p_{1}}b_{1}^{2}\succ_{p_{1}}b_{2}^{1}\succ_{p_{1}}b_{2}^{2}\succ_{p_{1}}b_{3}^{1}\succ_{p_{1}}b_{3}^{2}$
$c_{1}$ $:s_{1}^{1}\succ_{c_{1}}s_{2}^{1}\succ_{c_{1}}s_{3}^{1}\succ_{c_{1}}s_{1}^{3}\succ_{c_{1}}s_{2}^{3}\succ_{c_{1}}s_{3}^{3}\succ_{c_{1}}p_{1}$
$e$ $:s_{1}^{3}\succ_{e}s_{2}^{3}\succ_{e}s_{3}^{3}\succ_{e}q$
$f$ $:q\succ_{f}r$
Figure 8: Example of an Exact-MM instance, with $K=2$, and some of the preference lists in the transformed instance of ca-ml-lq-cl.
We now show that a maximal matching $M$ of size $K$ in $G$ implies a student-complete pair-stable matching $M^{\prime}$ in $I$ that respects lower quotas. For every pair $(u_{j},w_{i})\in M$, we add to $M^{\prime}$ either the pairs $(s_{i}^{1},c_{j})$ and $(s_{i}^{1},b_{i}^{1})$, or the pairs $(s_{i}^{2},c_{j})$ and $(s_{i}^{2},b_{i}^{2})$, depending on whether $s_{i}^{1}$ or $s_{i}^{2}$ has $c_{j}$ in her preference list. If $(s_{i}^{1},c_{j})\in M^{\prime}$, then we add $(s_{i}^{2},d_{i}^{1})$ to $M^{\prime}$, otherwise we add $(s_{i}^{1},d_{i}^{1})$. We also add $(s_{i}^{3},d_{i}^{2})$ and $(s_{i}^{4},d_{i}^{1})$ to $M^{\prime}$. If $w_{i}$ has no partner in $M$, we add the following pairs to $M^{\prime}$: $(s_{i}^{1},d_{i}^{1})$, $(s_{i}^{2},d_{i}^{2})$, $(s_{i}^{3},e)$, and $(s_{i}^{4},d_{i}^{1})$. Additionally, let $j_{k}$ ($1\leq k\leq K$) be a sequence such that $w_{j_{k}}$ is matched in $M$. If $(s_{i}^{1},c_{j_{k}})\in M^{\prime}$ for some $s_{i}^{1}$, add $(p_{j_{k}},b_{i}^{1})$ to $M^{\prime}$, otherwise add $(p_{j_{k}},b_{i}^{2})$ to $M^{\prime}$. Finally, add pairs $(q,f)$ and $(r,f)$ to $M^{\prime}$. We close every course in $B$ that has no students matched to it. The matching $M^{\prime}$ is student-complete, because for every student $s\in S$, $T(s)=O(C_{M^{\prime}}(s))$. It also respects the lower quotas of every course that appears in $M^{\prime}$. Course $f$ has two students, and every open course in $B$ and $D^{1}$ has two students too. We show below that it is pair-stable.
Lemma 9.
$M^{\prime}$ is a pair-stable matching.
Proof.
We prove the pair-size stability of $M^{\prime}$ by showing that no student can belong to a blocking pair. Every $s_{i}^{1}$ is matched with either some $c_{j}$ and $b_{i}^{1}$, which is her most preferred course, or with $d_{i}^{1}$. In the second case, $b_{i}^{1}$ is closed, so she cannot form a blocking pair with it. Similarly, every $s_{i}^{2}$ is matched with either some $c_{k}$ and $b_{i}^{2}$, her most preferred course, with $d_{i}^{1}$, or with $d_{i}^{2}$. If $s_{i}^{2}$ is matched with $b_{i}^{2}$ and $c_{k}$, or with $d_{i}^{1}$, the same reasoning applies, and if she is matched with $d_{i}^{2}$, then $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$ and $(s_{i}^{4},d_{i}^{1})\in M^{\prime}$, so there is no blocking pair containing $s_{i}^{2}$. The student $s_{i}^{4}$ is always matched with the only course in her preference list, $d_{i}^{1}$.
We now consider $s_{i}^{3}$. This student is matched with either $d_{i}^{2}$, her most preferred course, or with $e$. In the second case, $(s_{i}^{2},d_{i}^{2})\in M$ and, because $M$ is maximal in $G$, $u_{j}$ and $u_{k}$ are matched with some $w_{l}$ and $w_{m}$. Thus, there exist $s_{l}^{1}$ (or $s_{l}^{2}$) and $s_{m}^{1}$ (or $s_{m}^{2}$) such that $c_{j}$ and $c_{k}$ are matched with them. Hence, $s_{i}^{3}$ cannot form a blocking pair with any other course. Moving on to students in $P$, every $p_{k}\in P$ is matched with some $b_{i}^{l}\in B$, where $l=1$ or $l=2$. Let $b_{j}^{m}$ be an open course that $p_{k}$ prefers over $b_{i}^{l}$. Then $b_{j}^{m}$ is matched with both some $s_{i}^{1}$ or $s_{i}^{2}$ and with a $p_{k}^{\prime}$ such that $p_{k}^{\prime}\succ_{MLS}p_{k}$. Thus, $p_{k}$ cannot form a blocking pair with $b_{i}^{l}$. Finally, $r$ is matched with the only course in her preference list, $f$, and $q$ is matched with $f$, too. Even though $q$ prefers $e$ over $f$, $e$ has filled its capacity, and every student enrolled in $e$ belongs to $S^{3}$, all of which appear in the master list before $q$. Therefore, no student belongs to a blocking pair. ∎
Now, assume that there exists a student-complete pair-stable matching $M^{\prime}$ in $I$. We now show how to get a maximal matching $M$ of size $K$ in $G$. First, as $M^{\prime}$ is student-complete, $r$ is matched with $f$. The minimum quota of $f$ is $2$, so we have that $(q,f)\in M^{\prime}$ too. However, $q$ prefers $e$ over $f$. Thus, because $M^{\prime}$ is pair-stable, $e$ must have $n_{2}-K$ students assigned to it that are better than $q$. Hence, there are $n_{2}-K$ students from $S^{3}$ matched to $e$ in $M^{\prime}$. In order to preserve pair stability, for each $s_{i}^{3}$ matched to $e$, $M^{\prime}$ contains the pairs $(s_{i}^{1},d_{i}^{1})$ and $(s_{i}^{2},d_{i}^{2})$, as well as the pair $(s_{i}^{4},d_{i}^{1})$.
As $M^{\prime}$ is student-complete, every student in $P$ is matched to some $b_{i}^{l}$, where $l=1$ or $l=2$. In order to satisfy the lower quota of $b_{i}^{l}$ and make sure $M^{\prime}$ is pair-stable, $(s_{i}^{l},b_{i}^{l})\in M^{\prime}$ too. As $M^{\prime}$ is student-complete, $s_{i}^{l}$ is matched to some $c_{j}$ and $M^{\prime}$ includes the pair $(s_{i}^{4},d_{i}^{1})$. As the lower quota of $d_{i}^{1}$ is $2$, $M^{\prime}$ must include the pair $(s_{i}^{m},d_{i}^{l})$, where $m=1$ if $l=2$ and vice versa. This implies that $(s_{i}^{3},d_{i}^{2})\in M^{\prime}$. Hence, for all $1\leq i\leq n_{2}$, if $s_{i}^{1}$ is matched to some $c_{j}$ then $s_{i}^{2}$ is matched to $d_{i}^{1}$, and if $s_{i}^{2}$ is matched to some $c_{k}$ then $s_{i}^{1}$ is matched to $d_{i}^{1}$. Thus, we can build a matching $M$ in $G$ of size $K$ by taking every pair of the form $(s_{i}^{l},c_{j})\in M$, of which there are $K$, and adding the pair $(u_{j},w_{i})$ to $M$. We now show that $M$ is a maximal matching in $G$.
Lemma 10.
The matching $M$ in $G$, of size $K$, is maximal.
Proof.
Looking for a contradiction, assume that it is not, so that there exists unmatched $u_{j},w_{i}$ that are adjacent to each other. This implies that in $M^{\prime}$, $c_{j}$ is unmatched and $s_{i}^{3}$ is matched with $e$. However, $s_{i}^{3}$ prefers $c_{j}$ over $e$ and $c_{j}$ prefers $s_{i}^{3}$ over being unmatched. Thus, $(s_{i}^{3},c_{j})$ is a blocking pair, a contradiction to the stability of $M^{\prime}$. Therefore, $M$ is maximal. ∎
Hence, there exists a maximal matching of size $K$ in $G$ if and only if there exists a student-complete first-coalition-stable matching in $I$, respecting lower quotas of every open course. As the first problem is NP-complete, so is the second one. ∎
6 Conclusion
We have introduced a new model of the Course Allocation problem, which allows for courses with different numbers of credits and constraints such as courses that run at overlapping times. We have considered four natural definitions of stability in this context, and for each one we studied the problem of determining whether a matching is stable and finding a stable matching if one exists. Additionally, we studied the problem with master lists and lower quotas, and considered the problems of finding a stable matching and a maximum size stable matching.
The main problem remaining is finding a coalition-stable matching under arbitrary preferences, and determining whether one always exists. Another possible research direction is studying stability under additively separable preferences, where students assign a utility to a course and prefer subsets of courses where the sum of utilities is higher cechlarova_pareto_2018 . A restriction of additively separable preferences is lexicographic preferences, where students compare two bundles of courses by first checking the first course of each bundle, then the second course, and so on cechlarova_pareto_2018 . If we also require that students participating in a blocking (according to additively separable or lexicographic preferences) coalition do not reduce their total number of credits, then these definitions are more strict than first-coalition stability but less so than coalition stability.
Other possibilities include finding stable matchings with the lowest number of blocking pairs/coalitions when there is no stable matching, approximation algorithms for finding a maximum size stable matching, or modifying course upper quotas to ensure the existence of a stable matching, following the work of Nguyen and Vohra nguyen_2018_near .